A Complete Bibliography of Publications in Parallel Algorithms and Applications

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.03

Title word cross-reference

1 [AF93]. 1∗ [HSO96]. 2 [CLL95]. d [TJ03, Td96]. k [LQ96, SAOKM01]. L_{0} [GTA97]. m [SWT98]. n [SAOKM01]. O(log_{4} n) [DS04]. O(log p) [DD00, TFT02]. S_{n} [AR94].

-ary [SAOKM01, SWT98]. -Cubes [SAOKM01]. -D [CLL95]. -Dimensional [Td96]. -Neighbours [TJ03]. -Stable [GTA97].

2000 [VZ03]. 26 [Sto94].

5 [KCMC96].

Achieve [Rao96]. Action [CCZ95]. Activity [Cha98]. Adaptative [Ten96]. adaptive [SU03]. Admitting [AL97]. Advanced [Pla00]. Affine [LX02]. Age [BW95, KR96, KR97]. Aided [AY93]. Algebra [MAGE94]. Algebraic [Dje01]. Algorithm [AGGR95, ABC^{+}99, Als02, AR94, AS97, ABEA97, BEK97, BCC95, BA95, BX94, CS94, CS95, CJK93, DV97, DHM94, DS95, DS00, EY94b, FGS97b, FML95, GI96, GC94, GP94, HE96b, HLN95, JLT94, KR96, KR97, KJ95a, KJ95b, Kip96, KS95, LOSZ94, LO95, LY02, MW00, MB95, MM94, MM95, OSS^{+}93, OA94, Pap94, RDK00, SS94a, Stp03, SNS94, SWMC95, SWC99, Thu93, VMG03, YLE94, DS04, FM98, GSZ02, GSBM97, HZ04, LTCT97, OAE04, Pur03, Sog97].

Algorithmic [MSBK01]. Algorithms [AAD99, AAD00, AKT^{+}97, Ano94, ATG95a, ATG95b, AD94, Bah99, BBE95, Baz96, BV99b, BF94, Bha03, BC96, Bog94b, BL94b, Cal96, CLZ96, Cha98, CJP04, DFZ94, Dat01, DSVF95, DRS95, DKKS96, Dou96, ENSS93, EC95, GSA94, GA04, Guh98, HE96a, HKC95, HCL93, KCMC96, KRT98, Lan96, Li97, MKD97, MPS96, ND94, OBHS00, PB95, PLZ98, Pap95, SS95, Sto96, Sto94, TR02, TRFM95, TK95, BV98, BV99a, BT03, CE97, CM03, DRT94, EA97, MS00b, SU03, Ten96, UH97]. Alignment [ETN94]. All-Pairs [CS94]. Allen [DV97]. Allocation [ETN94, Zho94]. Alternate [Rao94]. Alternating [EJ94, Suk96, GSBM97]. Among [AS99]. Analog [DSVF95]. Analyser [MRS94]. Analysis [Cel93, CC02, DMSD95, FGS97a, HB97, KR97, KRT98, LPW97, NB99, OBHS00, SNS94, YXZ96, SAW02]. Analyzing [Cha98]. Angular [AR94]. Annealing [BC96, DLMS96, MKD97, YLE94]. Announcement [SS95]. Application [Als02, DZP00, LM98, SE95, ZZ04]. Application-Specific [LM98]. Applications [Ano94, BA95, Cel93, CPT00, CLL95, CC02, LCFL04, LO95, SS95, Sto94, GA04]. applied [GSBM97]. Approach [AR94, CE96, DKKS96, RTDL94, TK95, Zho94]. Approximate [Pin02, Spe94]. Approximation [Li97]. Aquifer [PS99]. Arbitration [MCMN95]. Architectural [SWC99]. Architecture [AY93, Bro97, CJP04, ENSS93, HNRW02, MRS94, SP95, WE93, HZ04]. Architectures [ABEA97, ES93, GP94, MS00a, PS95, DRT94, UH97]. Arcs [AB97]. Area [CW95]. arithmetic [BD95]. Array [AE97, Bro97, Cel93, CLL95, EW97, Fid01, LM98, MM04, MRS94, PMJ99a, PMJ99b, PLZ98, Pla00, SCM00, BT03, ME97].

Arrays [AF93, AKT^{+}97, BFM97, CES00, CLZ96, CT96, EC94, GE94, PA96, Tsa93]. Arrival [WE93]. Artificial [MAGE94]. ary [SAOKM01, SWT98]. Assignment [MB95, OA94, Pin97]. Asynchronous [Baz96, CD97, HNL97a, HNL97b, KCMC96, LN95, WBE94a, WBE94b]. Automata [LWC96]. Automatic [DZ96, SSBK99]. Axiomatic [SCG00]. Axiomatization [GC94].

B&B [GC94]. Background [SCM00]. Backpropagation [CE95b, SP95]. Backpropagation-Cauchy [CE95b]. Backward [AAD00]. balancer [MB04]. Balancing [EC95, GST95, Hen96, QA94, SEY97, Td96, MTR03]. Banded [RDK00]. Based [DS95, DS00, Ger03, KCMC96, LX02, MRS94, MM95, Qsn95a, Qsn95b, Skl02, SM01, WH96, MB04]. Bases [LL93]. Basic [UH97]. Behaviour [AL97]. Benchmarks [EK96]. Best [GC94]. Best-First [GC94]. Biharmonic [ME99]. Binary [HSO96, SKK04]. Bipartite [KJ95a, KJ95b, KS97]. Bit [Bha03, SSL04]. bit-serial [SSL04]. Bitonic [AB94]. Bits [FT96]. Blackboard [AY93]. Block [BV99a, BV99b, CS93, ES94, EB94b, GR94, GP94, MT95, ME99, Nik96, Sch95, XPJG94, Kon95, OV03]. Block-Finding [Nik96]. Block-Jacobi [BV99a, BV99b, OV03]. BMARS [BHW00]. Bordered [EB94b]. Both [PN96]. Boundary [AK95, KR95]. Boundary-Conforming [AK95]. Bounds [AKT^{+}97, CS93]. Bratislava [VZ03]. Breadth [LO95]. Breadth-First [LO95]. Bridges [ND94]. Broadcast [WL01]. Broadcasting [FGSxx, Sto96, Sto93]. BSP [SCG00]. BSP-Style [SCG00]. BSR [MS00b]. Bucket [Che97]. Buffered [PF01]. Buffers [KRT98]. Building [BE94, YXZ96]. Bulk [Cal96, CE97]. Bulk-synchronous [CE97]. Bus [FGS96, PLZ98, RS96, BT03, FGSxx].

Buses [Chu94, PA96, Li04].

Caching [Dou96]. Calculation [EY94b]. Calculations [Kha01]. Capacity [Dje01]. Capon [WE93]. Carlo [DKKS96, ESSU03, MAD94, MST94, WMF94, ZZ04]. Cascade [CCY95]. Case [GI96, QL94, Bha97]. Cauchy [CE95b]. Cellular [LWC96]. Center [AS97]. Centers [Cha96]. CFD [EK96]. CGLS [YL99]. chains [BD95]. Channel [MB95]. Chaotic [Bah99]. Character [ES96]. Checking [BV95]. chip [SSL04]. Cholesky [HZ04, PMJ99a]. Chromosome [BC96, Bha97]. Circuit [BAT99]. Circular [AB97]. Circulation [AS01]. Class [AF93, BEW97, DZE95, Fid01, WBE94a, WBE94b]. Classification [ET95]. Clipping [Chr95]. Cliques [AB97]. Clone [BC96]. Cluster [FW95, WMF94, AAL04]. Clustered [CE96]. Clustering [LPW97]. CM [KCMC96]. CM-5 [KCMC96]. Co [SCM00]. Co-Partitioning [SCM00]. Coarse [DD00]. Coarse-Grained [DD00]. Code [PB00, Pin97, PS99, Sto93]. Codes [MCMN95]. Coefficients [BW95]. Cographs [LOSZ94]. Collection [TR02]. Colouring [Bha03]. Combining [DLMS96, SCM00]. Common [LY02, LM98]. Communication [DD00, ERA94, SM01]. Communications [Bro97, LC97]. Comparative [AA96, PZ93, AAL04]. Comparison [DRS95, EA95]. Compile [FGS97a, PCJR97]. Compile-Time [FGS97a]. Compilers [SS95, SS97a, SS97b]. Complement [Sch95]. Complementarity [BEW97]. Complete [HSO96, OA95]. Complex [Pin02]. Complexity [HNL97b]. Composite [MW00, Thu93, Thu97]. Compositions [SE96]. compression [UH97]. Computation [AL97, BL94b, CW95, CCY95, CPT00, FPS94, Hor94, LMR96, MCSW97, Nak96, PAZN96, SS94b, Sto96, GSBM97].

Computational [Chr95, DSVF95].

Computations

[Bog94a, Dey94, MPS96, Tsa93, Li04, MB04]. compute [BV98]. Computer [AY93, CJK93, CE95b, FGS97b, HLM94, HNL97b, JS93, KJ95a, KJ95b, LE95, MPS96, ME99, Pin02, MT97]. Computer-Aided [AY93]. Computers [Chu94, DE95, DD00, DES97, EK96, EA95, MST94, RS96, Sto96, YL99, FM98, HNL97a]. Computing [AB97, BCC95, BL96, BW95, HNL97b, SJL^{+}94, SWMC95, SWC99, WMF94, Ano00, QPPJ03, Sog97]. Concurrent [AY93]. Conference [Ano93]. Confidence [HB97]. Conforming [AK95]. Congestion [HSO96]. Conjugate [EW97, GGM04, GSZ02]. Connected [AD94, ABEA97, CJK93, Chu94, FGS97b, KJ95a, KJ95b, Sto96, TOR93, Tsa93]. Connection [AB94]. connectivity [YM04]. conquer [GSZ02]. conscious [MB04]. Conservative [Cal96]. Considerations [SM01]. Considering [ERA94]. Consistent [PAZN96]. Constant [MB95, MS00b]. Constrained [Mit96]. Constraint [BV95]. Context [BBE95]. Context-Free [BBE95]. Contraharmonic [YE95a]. Convergence [HKC95, HNL97b, LN95, NB99, SWMC95]. Convex [CLZ96, KJ95a, KJ95b, Nak96]. CORBA [MTR03]. Coscheduling [PF01]. Cost [AA04, Chr95, LO95, VMG03]. Cost-effective [AA04]. Cost-Optimal [LO95, VMG03]. Coupled [EJ94, YE95b]. Covariance [BFM97]. Cover [LOSZ94]. CRAGE [EJ94]. Cray [BAT99, EK96]. Creation [Pin97]. Cross [TRFM95]. Cross-Platform [TRFM95]. crossed [YM04]. Cubes [SAOKM01, YM04]. Cubic [CFMP02]. Cut [CFMP02]. Cutpoints [ND94]. cycle [YM04]. Cycles [DK96].

D [AF93, CLL95]. DAGs [LPW97]. Dap [DOS97]. Data [DKKS96, DZ96, ETN94, GE96, MKD97, SP95, SSBK99, TE93, BHW00, PZ03]. Data-Driven [TE93]. Database [BV95]. DDM [KCMC96]. De-Coupled [YE95b]. Deadlock [ZPBD04]. Deadlock-free [ZPBD04]. Debugging [JDMW96]. Decision [AG96, DZP00]. Decomposition [EY94a, HE95, QL94, TOR93, Kon95]. Decompositions [PZ93]. Decoupling [ES94, Kip96]. Defects [ESSU03]. Definite [CE95a, EO97]. Delivery [AG96]. demand [FM98]. demand-driven [FM98]. Density [FPS94]. Dependence [AKT^{+}97, GE96, TK95]. Dependences [LX02, PC01, PCJR97]. Depth [AS01, DS95, DS00]. Depth-First [AS01]. Depth-First-Search [DS00]. Derivation [BB94, Meg93]. Deriving [LY02]. Design [EG97, MW00, Pla00, SM01]. Designs [AE97, MM04]. Detection [ABC^{+}99, PB00, RF00]. Determining [PB95]. Determinism [Rao96]. Deterministic [Rao96]. Developing [SCG00]. Devoted [Sto94]. Diagonal [CS93, EB94b]. Diary [Ano93]. difference [Pur03]. Different [OBHS00]. Differential [BGRD00, DKKS96, Sim94]. Diffusion [GTA97, NB99]. Dimension [DOS97]. Dimensional [GTA97, KS95, ME99, Td96]. Dimensions [Dou96]. Direct [RTDL94]. direction [Sog97]. Directions [WE93]. Directions-of-Arrival [WE93]. Discrete [BQOQO02, Cal96, CE96]. Discrimination [DOS97]. Distance [Nik96, Sog97]. Distances [TJ03]. Distributed [AFT00, ABC^{+}99, AS97, AS99, BV99b, CW95, CS94, CPT00, DbT01, Dje01, DZ96, EC95, ET95, GSA94, GGM04, GA04, GP94, HNL97b, KGB94, LL93, LMR96, MCMN95, MPS96, MW00, RS96, YL99, Zho94, BV99a, HNL97a]. Distributed-Memory [GSA94, KGB94]. Distribution [Lau02, Td96, AA04]. Divide [HLN95, GSZ02]. Divide-And-Inner [HLN95]. Doa [WE93]. Doing [Wil00].

Domain [DMSD95, HE95, QL94].

Dominance [Sto96]. Dominating [PB95]. Dominator [Hor94]. double [YM04]. double-vertex-cycle-connectivity [YM04]. Downdated [KC93]. DPART [DZ96]. Driven [TE93, FM98]. Dynamic [CE96, EC95, GST95, TE93, MB04, TFT02, ZPBD04].

economics [Ano00]. Edge [HSO96, PB95]. Edge-Congestion [HSO96]. Edge-Packing [PB95]. Editorial [Ano01, Eva93]. Effective [PC01, Zho94, AA04]. Efficient [AAD99, AKT^{+}97, AD94, ABEA97, DHM94, DFZ94, Dat01, Li97, LOSZ94, MS96, MRS94, MS00b, OA94, PB95, Pin97, Pla99, Qsn95a, SWT98, Sim94, SKK04]. Eigenpairs [Kha01]. Eigensolutions [MCSW97]. Eigenvalue [LE95, Pin02]. Eigenvalues [EY94b, BV98]. Eigenvectors [LX02, BV98]. Eigenvectors-Based [LX02]. Electromagnetic [AFT00]. Electronic [WH96]. Element [PS99]. Elimination [AAD00, BB94, CE95a, EA94, MM94, VMG03]. Elliptic [AA96, DKKS96, Gol00, HE96b, KCMC96]. Embeddings [HSO96]. Employing [SAOKM01]. Emulations [SWT98]. Enabled [LCFL04]. Enclosing [CJK93]. Ends [Rao96]. Engineering [AY93]. Enhanced [EY94c]. Entropy [Bog94b]. Enumerating [DK96]. Environment [TRFM95, MTR03, QPPJ03]. Environments [SWMC95, SWC99]. Epidemics [MSCSxx]. Equation [BB94, GTA97, Qsn95a, Qsn95b]. Equations [AF93, ATG95a, ATG95b, BQOQO02, BGRD00, CD97, DZE95, DKKS96, EC94, Fid01, Gol00, HE96b, KR96, KGB94, MM94, MM95, Sch95, Sim94, Meg93]. Eratosthenes [LQ96]. EREW [BX94, Kon00, LOSZ94, LO95]. Estimation [DOS97, HB97, MS00a, SCM00, WE93].

Estimator [BFM97]. Evaluation [HLN95, Kip96, Pap95, SP95, Stp03, QPPJ03, SU03]. Event [Cal96, CE96]. Eventual [Rao96]. Evolution [HKC95, HE95, YLE94]. Evolutionary [CPT00]. Evolving [PN96]. Exact [KGB94]. Exclusion [AS99]. Execution [AF93]. Expected [AF93]. Experience [BL94b]. Experiment [SJL^{+}94, ZP95, PZ03]. Expert [TE93]. Explicit [EJ94, GGM04, Suk96, YE95b, AAL04, OAE04, Pur03]. Explicitly [FGS97a]. Exploiting [AR94]. Exponentiation [CCY95]. Extended [HSO96, Sog97]. Extension [LQ96]. Extraction [DV97]. Extrapolation [YE95a]. Extremal [MCMN95].

Facility [PL94]. Factorisation [EA95, CE97]. Factorization [BL94a, MM95, PMJ99a, Rao94]. FAN [AGLP01]. Fast [DRT94, EG97, Li04, LY02, MS96, ND94, PLZ98, Stp03, FCF03]. Fastest [GE93]. Fault [BT03, SEY97]. Fault-Tolerant [SEY97]. feedback [TFT02]. FFT [AD94, SNS94]. Field [PAZN96]. finance [Ano00]. Financial [BGRD00, CPT00, DZP00]. Find [AS97]. Finding [Cha96, CJK93, ND94, Nik96, SS94b, TK95, BD95]. Finite [PS99]. First [AS01, DS95, DS00, GC94, LO95]. Fixed [HNL97b, KS95, LN95, PS95]. Fixed-Dimensional [KS95]. Fixed-Size [PS95]. Flame [DES97]. Flexible [FCF03, ME97]. floating [SSL04]. floating-point [SSL04]. Flow [CCZ95, PZ93, GSBM97]. Folding [GE94, ME95]. Foreword [Ola96, VZ03]. Formalism [LL93]. Forms [Pin97]. formulation [GSBM97]. Fortran [KMBK96]. four [OAE04]. Fourth [ATG95a]. Fourth-Order [ATG95a]. FPGA [HZ04, Skl02]. FPGA-Based [Skl02]. Fractal [DOS97, UH97]. Framework [Ran98, AGLP01].

Frameworks [SCG00]. Free

[BBE95, FML95, Pin94, ZPBD04]. frog [ESSU03]. Full [SCM00, VMG03]. function [GSBM97]. Functional [Dey94, DKKS96]. Functions [BEK97, FPS94, MSBK01].

Gauss [BE96]. Gaussian [AAD00, VMG03]. General [ME94a, PS95, Kon00]. Generalized [BL94a, CLL95, ME94b, Sto96, WL01]. Generating [SE96]. Generation [AK95, Gol00]. Genetic [BBE95, HKC95, MKD97, MW00, SSBK99, TRFM95]. Geometric [Als02, BF94]. Geometry [DSVF95]. Givens [Kon00, RTDL94]. Global [Bro97, DE95]. GMRES [SAW02]. Gossiping [FGS96, KRT98]. Gradient [EW97, GGM04, GSZ02]. Grained [DD00]. Gram [MM95, OBHS00]. Graph [AS97, Bha03, DLMS96, DK96, KJ95a, KJ95b]. Graph-Partitioning [DLMS96]. Graphs [Als02, CFMP02, Cha96, ERA94, KS97, ND94, OA95, PB95]. Gray [Sto93]. grey [SKK04]. Grid [AK95, Gol00, LCFL04]. Grid-Enabled [LCFL04]. Grids [HSO96, Ran98, Thu93, Thu97]. Grounding [AFT00]. Group [EJ94, Suk96, YE95b, AAL04, OAE04]. Growth [WMF94]. guided [TFT02]. GVT [BCC95].

Hamiltonicity [LOSZ94]. Hardware

[Skl02]. Hardware/Software [Skl02]. Hash [TR02]. Hashing [Pap94]. Hawk [HBT97]. Heap [DHM94]. Heat [LWC96]. Heavily [EC95]. held [VZ03]. Helpful [DLMS96]. Hessenberg [Qsn95a, Qsn95b]. Heterogeneous [CE96, CC02, ENSS93]. Heuristics [DRS95, LPW97]. HICSS [Sto94]. HICSS-26 [Sto94]. Hierarchical [SSBK99, Sog97]. High [Dje01, LCFL04, MST94, KMBK96]. High-Level [Dje01]. High-Performance [LCFL04]. Higher [MS00a, MSBK01].

Histogramming [JS93]. Homogeneous [EC95]. Householder [MM94]. HP [LCFL04]. Huard [MT97]. Hull [CLZ96, Nak96]. Hurricane [SJL^{+}94]. Hyperbolic [ATG95b]. Hypercube [AD94, Cel93, CS95, FML95, SS94a, Bha97]. Hypercube-Connected [AD94]. Hypercubes [SEY97, She00, Sto93, BV99a]. Hypercubic [BF94]. Hyperstar [AAD99].

i860 [LE95]. II [BV99b, PMJ99b]. Illustrated [SCM00]. Image [AE97, BET95, Chr95, ENSS93, ZP95, UH97]. Implementation [AE97, BAT99, BFM97, CE95b, DOS97, LE95, SP95, SWC99, WL01, WMF94, HZ04, TFT02]. Implementations [YE95b, SAW02]. Implicit [BE96, EA94]. Improved [Als02, DS00, FW95, PF01]. Incomplete [Sto93]. incompressible [GSBM97]. Incremental [Cha98]. Independent [CJP04]. Inference [HE96a]. InfiniBand^{T M} [ZPBD04]. Influences [SWC99]. Inhomogeneous [Ran98]. Inner [HLN95]. Integer [AGGR95, FGS97b, SE96]. integral [Ten96]. Integrals [EB94a]. Integration [EB94a, MTR03, ESSU03, SU03]. Interconnect [SM01]. Interconnection [FGS96, QA94, FGSxx]. Interface [Ano00]. Interior [HLM94]. Interlocking [EG97, GE96, Rao94]. International [VZ03]. Interpolation [SAOKM01]. Intertask [LC97]. Interval [May96, PB95]. Introduction [Lan96, SS97a, Sto94]. Inverse [Pin02]. Inversion [ME94a, MAD94]. IO [ZZ04]. Iso [Pla99]. Iso-Planes [Pla99]. Issue [GA04, Lan96, SS95, SS97a, Sto94]. Iteration [PS95, PCJR97]. Iterations [Dey94, LN95, ME95]. Iterative [Baz96, CD97, HNL97b, KM96, MAGE94, ME94b, MPS96, ME99, XPJG94, XPJG96, YE95b, AAL04, HNL97a].

Jacobi [BV99a, BV99b, OV03]. Janus [Sog97]. Java [LCFL04]. Join [TR02]. Jordan [BE96]. Journal [SS95, GA04].

Kennedy [DV97]. Kernel [SCM00]. Knapsack [AGGR95, Ten96, Meg93]. Knowledge [LL93]. Kohonen [BE94]. Kutta [YE95a].

Lagrange [SAOKM01]. Laminate [MW00]. Lanczos [GP94, MCSW97, Pin94, SWMC95, SWC99]. Lanczos-Type [Pin94]. Languages [BBE95, SS97a, SS95]. Large [BEW97, Dey94, HLM94, HE95, PMJ99a, PMJ99b, QL94, Pia03]. Large-Scale [Dey94]. Lattice [YXZ96]. Layered [TOR93]. LCS [MS00b]. Leaky [PS99]. leap [ESSU03]. leap-frog [ESSU03]. Least [YL99]. Leftmost [Pin02]. Level [AE97, Dje01, CM03, SKK04]. Library [Wil00]. Lightly [EC95]. Limit [HB97]. Limited [Dje01]. Line [JDMW96]. Linear [Bah99, BEW97, BE96, BQOQO02, BB94, Bro97, CLZ96, CE95a, CLL95, CD97, EC94, EA94, EB94b, EY94c, EW97, Fid01, GE96, HLM94, KCMC96, KGB94, KC93, KS95, MAGE94, MRS94, MT95, MM94, MM95, PMJ99a, PMJ99b, PLZ98, Rao94, RDK00, TK95, BT03, DS04, Kon00]. Liouville [FPS94]. Liquid [Hen96]. LIS [MS00b]. List [DD00]. Load [EC95, GST95, Hen96, Lau02, QA94, Td96, MTR03, MB04]. load-balancer [MB04]. Loaded [EC95]. Local [CW95]. locality [MB04]. locality-conscious [MB04]. Localization [BL96]. Locally [TOR93]. Location [PL94]. Logic [BB94, LL93]. Longest [LY02, LM98]. Loop [ERA95, PC01, PCJR97, TFT02]. Loops [DV97, EB95, LX02]. Low [AE97]. Low-Level [AE97]. Lower [AKT^{+}97].

Machine [AB94, BL94a, Cel93]. Machines [DZ96, ETN94, MSCSxx, SE95, TR02, TTS00]. Magic [EG97]. Management [SCM00]. Managing [DHM94]. Manipulation [Bha03]. Mapped [AF93]. Mapping [AKT^{+}97, ABEA97, CT96, CLL95, DRS95, EB95, MKD97, MPS96, Pla99, TOR93, Tsa93]. maps [Bha97]. Markov [BD95]. Massively [BC96, HLM94, KC93, SSL04, SKK04]. Matching [Als02, DS95, DS00, KJ95a, KJ95b, KS97, OA95, Spe94, DS04, ME97]. Matrices [BEK97, GR94, Kha01, Nik96, Pin02, BV98, EA97, Pia03]. Matrix [BQOQO02, EY94b, EA95, GSA94, GE93, KM96, ME94a, MAD94, MM95, PMJ99a, PMJ99b, PA96, Pia03, RTDL94, WBE94a, CE97, DRT94, Li04]. Max [CFMP02]. Maximal [DS95, DS00, DS04]. Maximum [AB97, OA95]. Mean [YE95a, BD95]. Means [Rao96]. Mechanism [WMF94]. Medians [Cha96]. Meiko [LE95]. Memory [AFT00, BV99b, DZ96, ETN94, GSA94, GP94, KGB94, MPS96, MS96, RS96, SCM00, SE95, TR02, YL99, BV99a, HZ04, MT97, OAE04]. Merging [BX94]. Mesh [ABEA97, CJP04, CJK93, Chu94, ES93, FGS97b, Guh98, JS93, KJ95a, KJ95b, MS96, MB95, Nak96, Sto96, Tsa93, MB04]. Mesh-Connected [ABEA97, CJK93, Chu94, FGS97b, KJ95a, KJ95b, Tsa93]. Mesh-Independent [CJP04]. Meshes [BV99b, OSS^{+}93, Sto96, Td96, OV03]. Message [Baz96, DRS95, Pap95, TOR93, QPPJ03]. Message-Passing [DRS95, TOR93, Pap95]. Method [AFT00, BW95, Baz96, CCZ95, CE95b, DE95, ES94, EJ94, EA94, EY94a, EY94c, EW97, GMP95, GGM04, GE96, HLM94, HE95, Hen96, HNL97b, KR95, Kha01, MAD94, MCSW97, QL94, Qsn95a, Qsn95b, Rao94, Sch95, SE96, Suk96, WE93, YE95a, YL99, BD95, HNL97a, MT97, Pur03]. Methodology [DZP00]. Methods [BEW97, CD97, DZE95, EA95, GTA97, MAGE94, ME94b, MT95, ME99, Sim94, SJL^{+}94, WBE94a, WBE94b, XPJG94, XPJG96, YE95b]. Metrics [LMR96]. Migration [MW00]. MIMD [Bha97, BL94b, DRT94, GSA94, MT97, UH97]. Minimum [LOSZ94]. Mining [BHW00]. Minitrack [Sto94]. Modal [LL93]. Model [AY93, Hen96, KC93, Pla99, RF00, Kon00, MS00b]. Modeling [CC02, GST95, MSCSxx, Skl02, AA04]. Modelling [CE96]. Models [LMR96, Tal01, TE93, YXZ96, AC99]. Modified [BFM97, MCSW97, MM95, OAE04]. Modular [CCY95, PS95]. Moments [AFT00, MS00a]. Monitoring [SWMC95]. Monotone [HE96b, HNL97b, HNL97a]. Monotone-Schwarz [HE96b]. Monte [DKKS96, ESSU03, MAD94, MST94, WMF94, ZZ04]. morphological [SKK04]. Motion [SCM00]. MPI [ZZ04]. MPI-IO [ZZ04]. Multi [AS01, BEW97, GTA97, YXZ96, CM03]. Multi-Dimensional [GTA97]. multi-level [CM03]. Multi-Parameter [BEW97]. Multi-Scalar [YXZ96]. Multi-Token [AS01]. Multicast [SEY97, SS94a]. Multicommodity [CCZ95, PZ93]. Multicomputers [AD94, BQOQO02, CS95, MKD97]. Multicriteria [DZP00]. Multidimensional [EB94a, QL94]. Multigrid [Dou96, DES97, Pap95]. Multilayer [ABEA97]. Multiple [BX94, Chu94, SP95, Sto96]. Multiplication [GSA94, GE93, PMJ99b, Pia03]. multiplications [DRT94]. Multiprocessor [GSA94, LC97, LPW97, HZ04]. Multiprocessors [AFT00, CD97, DRS95, KGB94, Pap95, SS94a, TOR93, OAE04]. Multiselection [She00]. Multisplitting [BEW97, DZE95, ME94b, PS95, WBE94a, WBE94b, XPJG96]. Multisplittings [Mit96, RM95]. Mutual [AS99, MS96].

natural [AA04]. Nearest [PLZ98]. Negligible [LC97]. negotiations [MB04]. Neighbor [PLZ98]. Neighboring [AS99]. Neighbours [TJ03]. Nested [DV97, EB95, LX02, MSBK01, PCJR97]. Nests [PC01]. Net [BE94]. Nets [ADVE97, Dje01, HE96a]. Network [AAD99, CCZ95, CE95b, MRS94, PL94, Pin94, QPPJ03, SP95, SNS94, CM03]. Networks [AK95, AAD00, AGGR95, AS01, BF94, CW95, Cha98, CC02, DbT01, ET95, FGS96, MAGE94, PN96, QA94, TOR93, ZP95, FGSxx, ZPBD04]. NEUCOMP2 [ES96]. Neural [ADVE97, BE94, CE95b, ET95, HE96a, MAGE94, PN96, SP95, TOR93]. Newton [HNL97a]. Node [TJ03]. Node-to-Node [TJ03]. Nodes [AS99]. Non [AF93, BL94b, CD97, Fid01, LC97, PC01, PCJR97, GSZ02]. Non-Linear [CD97]. Non-Negligible [LC97]. Non-Numeric [BL94b]. Non-Uniform [PC01, PCJR97]. Nonlinear [DZE95, Dey94, HE95, HE96b, ME99, PS99, QL94, WBE94b]. Normalized [GGM04]. NUMA [HNRW02, LC97]. number [DS04, MS00b]. Numeric [BL94b]. Numerical [AK95, BQOQO02, SU03]. Numerics [VZ03].

O.D.E. [YE95a]. Object [BV95, TR02]. Object-Oriented [BV95, TR02]. objects [HBT97]. On-Line [JDMW96]. One [ME99]. onto [AF93, Tsa93]. OOP [AY93]. Operations [KM96, PMJ99a, PMJ99b, PA96, Wil00, SKK04]. Operator [LN95]. Optical [PA96, Li04]. Optimal [BFM97, BF94, CES00, Che97, CJK93, DD00, ES93, ERA95, Guh98, HCL93, KJ95a, KJ95b, KS97, LO95, PL94, She00, Sto96, TK95, LTCT97, VMG03]. Optimality [DV97]. Optimising [SS95, SS97a]. Optimization [DE95, FGS97a, HKC95, Mit96, RM95, WH96, Ano00, CM03]. Optimizing [GSA94, HNRW02]. Optimum [SS94b]. Option [Ger03]. Order [ATG95a, ATG95b, MS00a, MSBK01, Sim94]. Ordered [LO95]. Ordering [BC96]. Orders [May96]. Ordinary [KC93, Sim94]. Oriented [BV95, TR02]. Orthogonal [EY94a, MM95]. Overlap [Che97]. Overlapped [AD94]. Overlapping [BCC95, QL94].

P.A.A. [Ano93]. Packing [PB95]. Pairs [CS94]. Pancake [QA94]. papers [VZ03]. Parabolic [ATG95a, KR96]. Paradigm [LWC96]. Paradigms [AL97]. Parallel [AA96, Akl93, AB97, AL97, AAD00, AG96, AGGR95, Als02, Ano94, ATG95a, ATG95b, BV98, Bah99, BEW97, BL96, BW95, BV95, BAT99, BB94, Bha03, BC96, Bha97, BET95, Bog94a, BA95, BL94b, BE94, BGRD00, Cal96, Cha96, Cha98, CE95a, CCY95, CJK93, CPT00, Chr95, CE95b, CM03, DHM94, DFZ94, DS95, DS00, Dat01, DE95, DSVF95, DD00, Dey94, DLMS96, DRS95, DKKS96, DES97, DZ96, ES93, ERA94, ETN94, EK96, ES94, EA94, EB94a, EB94b, EY94c, EA95, ET95, EB95, ES96, EO97, Eva97, FGS97a, FT96, FPS94, GSA94, GMP95, GC94, Ger03, Gol00, GA04, GP94, GTA97, HLM94, HB97, HC02, HE96b, Hor94, HCL93, HLN95, JDMW96, JLT94, KR95, KR96, KCMC96, KJ95a, KJ95b, KS97, Kip96, KC93, Kon00, KS95, LC97]. Parallel [Lan96, LL93, LE95, LMR96, Li97, LQ96, LWC96, MCMN95, MSCSxx, MS00a, ME94b, MPS96, May96, MT95, Mit96, ME99, MST94, MCSW97, MM94, MM95, ND94, Nik96, OBHS00, OA94, OAE04, PB00, PB95, Pap94, PAZN96, PZ93, PS99, Pin02, Qsn95a, Qsn95b, RTDL94, Rao94, RDK00, RM95, SAOKM01, Sch95, SE96, Sim94, SS94b, SS95, SS97a, SJL^{+}94, Spe94, Sto94, Stp03, SM01, SNS94, SWMC95, SWC99, Tal01, TR02, TE93, TRFM95, VZ03, WBE94a, WBE94b, WL01, WH96, YE95a, YL99, YE95b, ZZ04, AGLP01, Ano00, BHW00, CE97, DS04, ESSU03, FM98, GSZ02, GSBM97, Kon95, Li04, MS00b, Pia03, PZ03, Pur03, SSL04, SU03, SAW02, SKK04, TFT02, Ten96]. Parallelisation [LX02]. Parallelising [SS97b]. Parallelism [ADVE97, AR94, Bog94b, DV97, DMSD95, Bha97]. Parallelization [BEK97, CCZ95, PC01, FCF03]. Parameter [BEW97, KR96, Suk96]. Parlett [BEK97]. Parsing [BBE95]. Partial [BGRD00, MCSW97]. Particle [CJP04, WMF94, ZZ04]. Partition [GI96]. partitioned [HBT97]. Partitioning [Dat01, DLMS96, DZ96, ES94, OBHS00, PCJR97, Ran98, SCM00, SSBK99, Thu93, Thu97, LTCT97]. Parts [WH96]. Passing [Baz96, DRS95, TOR93, Pap95, QPPJ03]. Path [CS94, LOSZ94]. Pattern [DFZ94, SM01]. PC* [OA95]. PDE [AA96, HNRW02, KCMC96, MPS96, YXZ96]. Software [Skl02]. Perceptron [ABEA97]. Perfect [OA95]. Performance [CC02, FW95, HC02, HNRW02, KR97, KMBK96, LCFL04, LPW97, PAZN96, SP95, SNS94, QPPJ03]. Permutation [KS97]. Perpendicular [Bog94a]. Perturbation [KR97, Suk96]. Perturbed [Dey94, KR95]. Petri [Dje01]. Phases [DD00]. Physical [WH96, Bha97]. PIE [EA94]. Pipeline [AGGR95]. Pipelined [PLZ98, BT03, Li04]. Places [Dje01]. Planar [CLZ96, DK96, Gol00]. Plane [OA94]. Planes [Pla99]. Platform [TRFM95]. Point [HLM94, HNL97b, LN95, SSL04]. Points [Nak96, OAE04]. Polygons [Guh98, HCL93]. Polyhedral [Wil00]. Polynomial [GR94, HLN95, Kip96, Stp03]. Polytope [Pla99, RF00]. Portable [CJP04]. Porting [CC02]. Positive [CE95a, EO97]. Potential [Bog94b]. power [AC99]. PPTran [KMBK96]. Practical [FGS97b].

PRAM [AC99, BX94, GI96, Kon00].

Preconditioned

[EW97, GGM04, XPJG94, XPJG96, YL99]. Preconditioner [GR94]. Preconditioning [CS93, Pin02]. Prediction [HC02, PAZN96, SJL^{+}94]. Prefix [ES93]. Preprocessing [HCL93]. Price [Ger03]. Probabilistic [Rao96]. Problem [AGGR95, ABC^{+}99, Als02, CS94, DLMS96, GI96, Lau02, Li97, LY02, LM98, MB95, OA94, OA95, CM03, LTCT97, Pia03, Ten96, Meg93]. Problems [BEW97, CCZ95, Dou96, DES97, DZP00, ES96, GMP95, Guh98, HLM94, KR95, KR97, LE95, LWC96, ME99, PZ93, Sto96, Suk96, YL99, MS00b]. Processing [AG96, AE97, Chr95, Bha97]. Processor [AK95, AKT^{+}97, CES00, Cel93, CLL95, ERA95, GE94, Lau02, ME97, OV03, SKK04]. Processor-Time-Optimal [CES00]. Processors [ERA94, LM98, Pap94, Tsa93, DS04]. Product [AAD00, HLN95]. Program [FW95, LC97, SP95, Zho94]. Programming [EB95, HLM94, KS95, LCFL04, Tal01, TRFM95, AGLP01]. Programs [FGS97a, HC02, JDMW96, SCG00]. Properties [GE96, MKD97]. Property [EG97]. Protocol [AS99]. Prototyping [WH96]. Proximity [Guh98]. Pseudospectral [BGRD00]. Pvm [SE95].

QR [CE97, EA95, Kon95]. Quadrant [Rao94]. Quadratic [Kha01]. quasi [ESSU03]. quasi-Monte [ESSU03]. Quasilinear [GMP95]. Queueing [CE96, GST95]. QZ [CE97, EY94b, EY94a, EA95].

Random [DE95, FT96]. Randomized [DD00, Li97]. Range [Dat01]. Ranking [DD00]. RAP [Hor94]. RCC [VMG03]. Real [BFM97, EY94b, MS00a, Pin02, RTDL94]. Real-Time [BFM97, MS00a]. Recognition [DFZ94, ES96]. Reconfigurable [BL94a, JS93, KM96, MS96, MB95, Nak96, OSS^{+}93, PLZ98, PA96, Td96, BT03, Li04]. reconfiguration [ZPBD04]. Reconstruction [BC96, Bha97, PZ03]. Recurrence [AF93, Fid01, FML95, Meg93]. Recurrence-Free [FML95]. Recursive [ES94, Eva97]. Recycling [FT96]. Reduced [EJ94]. Reduction [EY94c]. Reductions [MM94, Pla99]. Regular [AF93, CT96, CC02, Pla00, SCM00, TTS00, TJ03]. Regularizing [MM04]. Related [Sto96]. Relaxation [WBE94a, WBE94b]. Relaxed [BEW97, DZE95]. Reliable [CS95]. Representations [Che97]. Resource [LMR96, PF01, AA04]. Restarting [SWC99]. Restoration [BET95, ZP95]. Restricted [KRT98]. Rigorous [DMSD95]. Ring [Lau02]. rings [BV99a]. Romberg [EB94a]. Routing [Sto93]. Row [CCZ95]. Row-Action [CCZ95]. Runge [YE95a]. runtime [HBT97].

s [HNL97b]. Salesman [Li97]. Scalability [SAW02]. scalable [Li04]. Scalar [YXZ96]. Scale [Dey94, HLM94, HE95, QL94]. Scans [RF00]. Scheduled [TE93]. Schedules [TK95]. Scheduling [AG96, BA95, ERA94, ERA95, LC97, LPW97, May96, TFT02]. Scheme [BE96]. Schemes [OBHS00, Pin94]. Schmidt [MM95, OBHS00]. Schur [Qsn95a, Qsn95b, Sch95]. Schwarz [GSBM97, HE96b]. Search [DS95, DS00, DE95, NB99, SSBK99]. Searching [BX94, Dat01]. Second [ATG95b, Sim94]. Second-Order [ATG95b]. Segmented [Chu94]. Segments [HCL93]. Selected [VZ03]. Selection [OSS^{+}93, QA94, RS96]. selections [MS00b]. Self [AS97, AS99, AS01, CS94, DbT01, HNRW02, MS96, PAZN96]. Self-Optimizing [HNRW02]. Self-Stabilizing [AS97, AS99, AS01, CS94, DbT01].

September [VZ03]. Sequences [LM98, Kon00]. Serial [DES97, SSL04]. Service [AG96]. Set [GI96, LTCT97]. Sets [DLMS96, PB95]. Shared [MS96, SE95, TR02, HZ04, MT97, OAE04]. Shared-Memory [TR02, HZ04]. Sheet [DES97]. Shortest [CS94]. Sided [MM94]. Sieve [LQ96]. SIMD [BL94b, ETN94, FM98, HLM94, KC93]. Simple [Guh98, OSS^{+}93]. Simulated [BC96, DLMS96, MKD97]. Simulation [BAT99, Cal96, Dje01, ES96, GST95, HNL97b, WMF94, ZZ04, FM98]. Simulations [MS96, MST94]. Simulator [BE94]. Single [Pin97, SP95]. Singular [Bah99, KR97, Suk96]. Singularly [KR95]. Size [PS95]. Skeletal [HC02]. skeleton [AGLP01]. Skeletons [MSBK01]. Small [KR96, Pia03]. Smallest [CJK93]. Software [CJP04]. Solution [AA96, BE96, BQOQO02, BGRD00, CD97, DES97, EA94, EB94b, Eva97, KGB94, LWC96, ME94a, MM94, MM95, Rao94, Sch95, YE95a, EO97, EA97]. Solver [Qsn95a, Qsn95b, OV03]. Solvers [BB94, Gol00, HNRW02, PMJ99b, AAL04]. Solving [ABC^{+}99, EY94c, HLM94, KR95, KCMC96, KC93, Li97, MT95, YL99, Kon00]. some [AC99]. Sort [AB94, TR02]. Sort-Hash [TR02]. Sorted [Nak96]. Sorting [CS95, Che97, Chu94, DbT01, DZP00, FGS97b, QA94, RS96]. Source [KMBK96]. Spaces [PCJR97]. Sparse [BEW97, PMJ99a, PMJ99b, Pin02, YL99]. Spatial [MSCSxx]. Special [GI96, GA04, Lan96, Sim94, SS97a, SS95, Sto94]. Specific [LM98]. SPECT [PZ03]. Spectral [BFM97, FPS94]. Speed [MCMN95, MST94, Pur03]. Speed-Up [MCMN95, Pur03]. Speedup [Cel93]. Splitting [LN95]. SPMD [SP95]. Squares [YL99]. Stabilizing [AS97, AS99, AS01, CS94, DbT01]. Stable [BQOQO02, GTA97]. Star [QA94]. State [HB97, TTS00]. Static [HC02, LPW97]. stationary [BD95]. Statistical [MS00a]. statistics [Ano00]. Stochastic [NB99]. Strassen [FML95]. Strategies [AA96, CCZ95, Thu97, Kon95]. Strategy [BL96, BB94]. stream [GSBM97]. Stride [EY94c]. Striding [Eva97]. string [ME97]. Structure [PB00]. Structured [AS99, PL94, Ran98, Tsa93]. Structures [MW00]. Study [AA96, BC96, DRS95, PZ93, AAL04, Bha97]. Sturm [FPS94]. Style [SCG00]. Subdefinite [Kha01]. Subdomains [QL94]. Subsequence [LY02, LM98]. Substitution [AAD00]. Supercomputers [BAT99]. Superunitary [AL97]. Support [AG96, LCFL04]. Surface [JLT94]. SVD [BV99a, BV99b, OV03]. Sylvester [Qsn95a, Qsn95b]. Symmetric [CE95a, HE95, Kha01, LE95, RTDL94, BV98, EO97, GSZ02]. Synchronous [Cal96, CE97]. Synergy [Akl93]. Synthesis [MS00a]. System [AG96, AS99, BV95, BB94, CE95b, DZE95, DZ96, LPW97, ME94a, PMJ99b, PLZ98, RDK00, BT03, HBT97, SSL04]. Systematic [MS00a]. Systems [AFT00, Bah99, BE96, Baz96, BV99b, CE96, CE95a, CD97, EC94, EA94, EB94b, EY94c, EC95, Eva97, HE95, KC93, LC97, MT95, PS99, QL94, Rao94, Sch95, SS97b, Skl02, TE93, AA04, BV99a, EO97, GSZ02, SU03]. Systolic [AE97, AKT^{+}97, BFM97, CES00, EC94, EW97, EG97, EA97, GE94, HE96a, KM96, LY02, ME95, ME94a, MAD94, WE93, LTCT97, OV03].

TAGE [Suk96]. Task [ERA94, LPW97]. Tasks [LC97]. taxation [BHW00]. technique [ESSU03]. Techniques [BET95, ME95, PC01, SWC99]. Terminating [Baz96]. Termination [ABC^{+}99]. test [Sog97]. Testing [HCL93]. Texture [DOS97]. th [LQ96]. Their [SWC99, YE95b]. Theory [GST95]. Three [EY94c, LM98]. Time [AF93, AKT^{+}97, BCC95, BFM97, BF94, CES00, DMSD95, FGS97a, MS00a, MB95, PCJR97, Sto96, DS04]. Time-Optimal [BF94, Sto96]. Timespace [CT96]. Toeplitz [BV98, EC94, EO97, EA97, GSZ02]. Token [AS01, Td96]. Tolerant [SEY97, BT03]. Tools [MM04]. Topologies [TJ03]. Topology [PN96]. Torus [KM96]. Track [SJL^{+}94]. Tracking [CJP04]. Traffic [SEY97]. Traffic-Balancing [SEY97]. Training [CE95b]. Transfer [LWC96]. Transformation [RTDL94, AGLP01]. Transformations [GE94]. transforms [FCF03]. Transients [AFT00]. Translator [KMBK96]. Transport [AR94, ZZ04]. Transpose [Pin94]. Transpose-Free [Pin94]. Transputer [AE97, Pin94, SNS94, ZP95]. Trapezoid [Als02]. Traveling [Li97]. Tree [AS97, AS99, AS01, DbT01, PL94, Tsa93, Ger03]. Tree-Structured [Tsa93]. Trees [HSO96, LO95, SWT98]. Trends [Tal01]. Triangle [CJK93]. Triangular [BEK97, PMJ99b, Qsn95b]. Triangulation [JLT94]. Tridiagonal [EY94c, Eva97, GR94, Sch95, GSZ02]. Tridiagonalizing [RTDL94]. Trinomial [Ger03]. Trinomial-tree [Ger03]. Two [Dou96, ERA95, MM94, Suk96]. Two-Processor [ERA95]. Two-Sided [MM94]. Type [Pin94].

Uncertain [BBE95]. Understanding [ENSS93]. Undirected [ND94]. Unified [AAD00]. Uniform [AF93, AKT^{+}97, Fid01, PC01, PCJR97, TK95, Meg93]. unit [SSL04]. Universal [EB95]. Unrolling [ERA95]. Unstructured [Pin97, MB04]. Updated [KC93]. Updating [BET95, Kon95]. Use [YXZ96]. Using [AY93, Bha03, ES96, FW95, GST95, Kha01, Li97, MAD94, MB95, MCSW97, MM94, Nik96, PA96, Pin02, Pla99, RTDL94, Rao96, SJL^{+}94, ZP95, DS04, ESSU03, GSZ02, HLM94, MS00b]. Utilization [PF01]. Utilizing [AG96].

Valuations [Ger03]. Value [KR95]. Variable [BW95]. Variant [FML95]. Vector [BAT99, GE93, MST94, BD95]. Vectorizable [XPJG94]. vectors [Sog97]. version [Sog97]. Versus [BL94b]. vertex [YM04]. Vertices [Bha03]. Via [BC96]. Virtual [SE95]. Visibility [Guh98, HCL93]. Vision [Cel93]. VLSI [GE94, ME97, PS95]. vorticity [GSBM97].

Warp [BCC95]. Wavefront [SS94b]. wavelet [FCF03]. Weak [HCL93]. Weight [OA95]. Weighted [OA95]. Weights [PN96]. Window [BCC95]. Within [BB94]. Workload [Ran98]. Workshop [VZ03]. Workstation [CC02]. Workstations [CW95, FW95, AAL04]. WZ [BL94a].

X [SWT98]. X-Trees [SWT98].

References

Abdullah:1996:CSP

[AA96] Abdul Rahman Abdullah and Norhashidah Hj. Mohd. Ali. A comparative study of parallel strategies for the solution of elliptic PDE’s. Parallel Algorithms and Applications, 10(1–2):93–103, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776220900.

Al-Ayyoub:2004:CEM

[AA04] Abdel-Elah Al-Ayyoub. Costeffective modeling for natural resource distribution systems. Parallel Algorithms and Applications, 19(4):257–265, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a714592660.

Al-Ayyoub:1999:EAH

[AAD99] Abdel-Elah Al-Ayyoub and Khaled Day. Efficient algorithms on the hyperstar network. Parallel Algorithms and Applications, 14(1):79–88, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777980766.

Al-Ayyoub:2000:UPA

[AAD00] Abdel-Elah Al-Ayyoub and Khaled Day. Unified parallel algorithms for Gaussian elimination with backward substitution on product networks. Parallel Algorithms and Applications, 14(4):253–269, April 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778714133.

Ali:2004:CSE

[AAL04] Norhashidah Hj. Mohd. Ali, Rosni Abdullah, and Kok Jun Lee. A comparative study of explicit group iterative solvers on a cluster of workstations. Parallel Algorithms and Applications, 19(4):237–255, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a714592659.

Ayoubi:1994:BSC

[AB94] R. A. Ayoubi and M. A. Bayoumi. Bitonic sort on the connection machine. Parallel Algorithms and Applications, 3(1–2):151–161, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779202349.

Akl:1997:CMC

[AB97] Selim G. Akl and Binay K.

Bhattacharya. Computing maximum cliques of circular arcs in parallel. Parallel Algorithms and Applications, 12(4):305–320, April 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776016229.

Aloisio:1999:ASD

[ABC^{+}99] G. Aloisio, P. Beraldi, M. Cafaro, F. Guerriero, and R. Musmanno. An algorithm for solving the distributed termination detection problem. Parallel Algorithms and Applications, 14(2):149–164, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777957924.

Ayoubi:1997:EMA

[ABEA97] R. Ayoubi, M. Bayoumi, A. Elchouemi, and B. Alhalabi. An efficient mapping algorithm of multilayer perceptron on meshconnected architectures. Parallel Algorithms and Applications, 11(3–4):273–285, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772742294.

Akl:1999:PSP

[AC99] Selim G. Akl and Lin Chen. On the power of some PRAM models. Parallel Algorithms and Applications, 13(4):307–319, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776633632.

Aykanat:1994:EOF

[AD94] Cevdet Aykanat and Argun Dervi. Efficient overlapped FFT algorithms for hypercubeconnected multicomputers. Parallel Algorithms and Applications, 4(1–2):91–110, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745102.

Al-Dabass:1997:PNN

[ADVE97] D. Al-Dabass, P. Vindlacheruvu, and D. J. Evans. Parallelism in neural nets. Parallel Algorithms and Applications, 11(3–4):169–185, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772742289.

Amin:1997:TIL

[AE97] S. A. Amin and D. J. Evans. Transputer implementation for low-level image processing systolic array designs. Parallel Algorithms and Applications, 10(3–4):283–299, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398800.

Aleksandrov:1993:EET

[AF93] Vassil Aleksandrov and Stefka Fidanova. On the expected execution time for a class of non uniform recurrence equations mapped onto 1D regular arrays. Parallel Algorithms and Applications, 1(4):303–314, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397701.

Ala:2000:MME

[AFT00] G. Ala, E. Francomano, and A. Tortorici. The method of moments for electromagnetic transients in grounding systems on distributed memory multiprocessors. Parallel Algorithms and Applications, 14(3):213–233, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778235468.

Alaghband:1996:UPP

[AG96] Gita Alaghband and Melodi Mosley Gates. Utilizing parallel processing: a decision support system for service delivery scheduling. Parallel Algorithms and Applications, 10(1–2):1–20, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776220892.

Almeida:1995:PAI

[AGGR95] Francisco Almeida, Felix García, Daniel Gonzalez, and Casiano Rodríguez. A parallel algorithm for the integer knapsack problem for pipeline networks. Parallel Algorithms and Applications, 6(2–3):87–91, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398619.

Aldinucci:2001:TPP

[AGLP01] M. Aldinucci, S. Gorlatch, C. Lengauer, and S. Pelagatti. Towards parallel programming by transformation: the FAN skeleton framework. Parallel Algorithms and Applications, 16(2–3):87–121, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777319791.

Aden:1995:BCN

[AK95] Jörg-Udo Aden and Norbert Köckler. Boundary-conforming numerical grid generation on processor networks. Parallel Algorithms and Applications, 5(3–4):251–268, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772400095.

Akl:1993:PS

[Akl93] Selim G. Akl. Parallel synergy. Parallel Algorithms and Applications, 1(1):3–9, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772399753.

Andronikos:1997:LTP

[AKT^{+}97] T. Andronikos, N. Koziris, Z. Tsiatsoulis, G. Papakonstantinou, and P. Tsanakas. Lower time and processor bounds for efficient mapping of uniform dependence algorithms into systolic arrays. Parallel Algorithms and Applications, 10(3–4):177–193, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http:

//www.dsclab.ece.ntua.gr/
~nkoziris/papers/jpaa-97-
abstract.html;   http://www.
informaworld.com/smpp/content~
content=a772398782.

Akl:1997:PAS

[AL97] S. G. Akl and L. Fava Lindon. Paradigms admitting superunitary behaviour in parallel computation. Parallel Algorithms and Applications, 11(1–2):129–153, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742766.

Alsuwaiyel:2002:IPA

[Als02] Muhammad H. Alsuwaiyel. An improved parallel algorithm for a geometric matching problem with application to trapezoid graphs. Parallel Algorithms and Applications, 17(2):147–155, ???? 2002. CODEN PAAPEC.

ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777924085.

Anonymous:1993:PCD

[Ano93] Anonymous. P.A.A. Conference diary. Parallel Algorithms and Applications, 1(3):253, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398096.

Anonymous:1994:PAA

[Ano94]  Anonymous.  Parallel algo-

rithms and applications. Parallel Algorithms and Applications, 3(1–2):163–164, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779202350.

Anonymous:2000:IPC

[Ano00] Anonymous. Interface of parallel computing and statistics, economics, finance and optimization. Parallel Algorithms and Applications, 15(1–2):1–2, 2000. CODEN PAAPEC. ISSN 1063-7192.

Anonymous:2001:E

[Ano01] Anonymous. Editorial. Parallel Algorithms and Applications, 16(2–3):83–85, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777319790.

Altekar:1994:EPT

[AR94] S. D. Altekar and A. K. Ray. Exploiting parallelism in a S_{n} transport algorithm: An angular approach. Parallel Algorithms and Applications, 4 (3–4):317–328, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314740.

Antonoiu:1997:SSD

[AS97] Gheorghe Antonoiu and Pradip K.

Srimani. A self-stabilizing distributed algorithm to find the center of a tree graph. Parallel Algorithms and Applications, 10(3–4):237–248, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398797.

Antonoiu:1999:SSP

[AS99] Gheorghe Antonoiu and Pradip K.

Srimani. Self-stabilizing protocol for mutual exclusion among neighboring nodes in a tree structured distributed system. Parallel Algorithms and Applications, 14(1):1–18, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777980762.

Antonoiu:2001:SSD

[AS01] Gheorghe Antonoiu and Pradip K.

Srimani. Self-stabilizing depthfirst multi-token circulation in tree networks. Parallel Algorithms and Applications, 16(1):17–35, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777308745.

Arigu:1995:PAF

[ATG95a] M. A. Arigu, E. H. Twizell, and A. B. Gumel. Parallel algorithms for fourth-order parabolic equations. Parallel Algorithms and Applications, 6(4):273–286, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772403914.

Arigu:1995:PAS

[ATG95b] M. A. Arigu, E. H. Twizell, and A. B. Gumel. Parallel algorithms for second-order hyperbolic equations. Parallel Algorithms and Applications, 5(1–2):119–128, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743845.

Adeli:1993:COM

[AY93] H. Adeli and G. Yu. A concurrent OOP model for computeraided engineering using blackboard architecture. Parallel Algorithms and Applications, 1(4):315–337, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772397702.

Bozyigit:1995:PSA

[BA95] Muslim Bozyigit and Amin A. Abdulghani. A parallel scheduling algorithm for parallel applications. Parallel Algorithms and Applications, 6(4):303–316, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772403916.

Bahi:1999:PCA

[Bah99] J. Bahi. Parallel chaotic algorithms for singular linear systems. Parallel Algorithms and Applications, 14(1):19–35, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777980763.

Bataineh:1999:PVI

[BAT99] Abdulla Bataineh, Mike Aamodt, and Kevin Thomas. A parallel and vector implementation of circuit simulation on Cray supercomputers. Parallel Algorithms and Applications, 14(2):109–118, March 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777957922.

Baz:1996:MTA

[Baz96] Didier El Baz. A method of terminating asynchronous iterative algorithms on message passing systems. Parallel Algorithms and Applications, 9(1–2):153–158, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772744477.

Berzigyarov:1994:LDP

[BB94] Parvaz K. Berzigyarov and Yuri P. Boglaev. Logic derivation of parallel linear equation system solvers within elimination strategy. Parallel Algorithms and Applications, 4(1–2):21–40, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745097.

Barsan:1995:PUC

[BBE95] L. Barsan, C. Barsan, and D. J. Evans. Parsing uncertain context-free languages by genetic algorithms. Parallel Algorithms and Applications, 7(3–4):209–228, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772402407.

Bhandarkar:1996:SMP

[BC96] Suchendra M. Bhandarkar and Sridhar Chirravuri. A study of massively parallel simulated annealing algorithms for chromosome reconstruction via clone ordering. Parallel Algorithms and Applications, 9(1–2):67–89, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772744468.

Baldwin:1995:OWA

[BCC95] Reid Baldwin, Moon Jung Chung, and Yunmo Chung. Overlapping window algorithm for computing GVT in time warp. Parallel Algorithms and Applications, 6(2–3):93–110, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398620.

Benzi:1995:AMM

[BD95] Michele Benzi and Tuğrul Dayar.

The arithmetic mean method for finding the stationary vector of Markov chains. Parallel Algorithms and Applications, 6(1):25–37, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397972.

Buhusi:1994:BKN

[BE94] C. V. Buhusi and D. J. Evans.

On building a Kohonen neural net parallel simulator. Parallel Algorithms and Applications, 2(3):229–244, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401610.

Barulli:1996:IGJ

[BE96] M. Barulli and D. J. Evans.

Implicit Gauss–Jordan scheme for the solution of linear systems. Parallel Algorithms and Applications, 10(1–2):145–159, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776220904.

Bakkaloglu:1997:PPA

[BEK97] B. Bakkaloğlu, K. Erciyeş, and Ç. K. Koç. A parallelization of Parlett’s algorithm for functions of triangular matrices. Parallel Algorithms and Applications, 11(1–2):61–69, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772742762.

Blaznik:1995:PUT

[BET95] P. Blaznik, D. J. Evans, and J. Tasič. Parallel updating techniques in image restoration. Parallel Algorithms and Applications, 7(3–4):237–253, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772402449.

Bai:1997:CMP

[BEW97] Zhong-Zhi Bai, D. J. Evans, and De-Ren Wang. A class of multi-parameter relaxed parallel multisplitting methods for large sparse linear complementarity problems. Parallel Algorithms and Applications, 11(1–2):113–127, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772742765.

Berthome:1994:TOG

[BF94] Pascal Berthomé and Afonso Ferreira. Time-optimal geometric algorithms in hypercubic networks. Parallel Algorithms and Applications, 4(3–4):169–181, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777314729.

Bellis:1997:OSA

[BFM97] Stephen J. Bellis, Peter J. Fish, and William P. Marnane. Optimal systolic arrays for realtime implementation of the modified covariance spectral estimator. Parallel Algorithms and Applications, 11(1–2):71–96, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742763.

Bunnin:2000:PPS

[BGRD00] F. O. Bunnin, Y. (Yi Ke) Guo, Y. (Yuhe) Ren, and J. Darlington. Parallel pseudospectral solution of financial partial differential equations. Parallel Algorithms and Applications, 15(1–2):3–13, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776636254.

Interface of parallel computing and statistics, economics, finance and optimization.

Bhandarkar:1997:PPC

[Bha97] Suchendra M. Bhandarkar. Parallel processing for chromosome reconstruction from physical maps — a case study of MIMD parallelism on the hypercube. Parallel Algorithms and Applications, 12(4):231–252, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776016214.

Bhalla:2003:PAC

[Bha03] Neelam Bhalla. Parallel algorithms for colouring vertices of graph using bit manipulation. Parallel Algorithms and Applications, 18(4):197–208, December 2003. CODEN PAAPEC.

ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713646633.

Bakin:2000:MTD

[BHW00] Sergey Bakin, Markus Hegland, and Graham Williams. Mining taxation data with parallel BMARS. Parallel Algorithms and Applications, 15(1–2):37–55, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776636256.

Interface of parallel computing and statistics, economics, finance and optimization.

Benaini:1994:GWF

[BL94a] A. Benaini and D. Laiymani. Generalized WZ factorization on a reconfigurable machine. Parallel Algorithms and Applications, 3(3–4):261–269, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779167061.

Breshears:1994:MVS

[BL94b] Clay P. Breshears and Michael A. Langston. MIMD versus SIMD computation: Experience with non-numeric parallel algorithms. Parallel Algorithms and Applications, 2(1–2):123–138, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745446.

See also the authors’ paper in Proceeding of the Twenty-Sixth Hawaii International Conference on System Sciences (IEEE Catalog Number 93TH0501-7), IEEE, p. xvi + 895 + xiv + 691 + xii + 654 + xv + 889, pp. 298–307, vol. 2 (1993).

Bao-Lin:1996:LSP

[BL96] Zhang Bao-Lin. Localization strategy for parallel computing. Parallel Algorithms and Applications, 9(3–4):185–193, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778707977.

Boglaev:1994:PPC

[Bog94a] Yuri P. Boglaev. On the parallel and perpendicular computations. Parallel Algorithms and Applications, 3(1–2):89–107, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779202305.

Boglaev:1994:EAP

[Bog94b] Yuri R. Boglaev. Entropy of algorithms and potential parallelism. Parallel Algorithms and Applications, 3(1–2):109–133, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779202306.

Benner:2002:NSD

[BQOQO02] Peter Benner, Enrique S.

Quintana-Ortí, and Gregorio Quintana-Ortí. Numerical solution of discrete stable linear matrix equations on multicomputers. Parallel Algorithms and Applications, 17(2):127–146, ???? 2002. CODEN PAAPEC.

ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777924084.

Broggi:1997:GCL

[Bro97] Alberto Broggi. Global communications on a linear array architecture. Parallel Algorithms and Applications, 11(1–2):27–43, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772742760.

Bourgeois:2003:FTA

[BT03] Anu G. Bourgeois and Jerry L. Trahan. Fault tolerant algorithms for a linear array with a reconfigurable pipelined bus system. Parallel Algorithms and Applications, 18 (3):139–153, September 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713812563.

Bassiliades:1995:CCP

[BV95] N. Bassiliades and I. Vlahavas. Constraint checking in a parallel object-oriented database system. Parallel Algorithms and Applications, 5(1– 2):129–147, ???? 1995. CO-DEN PAAPEC. ISSN 1063-

                                        7192. URL http://www.csd.
                                        auth.gr/~plk/publications/
                                        PAA-5.html;   http://www.
                                        informaworld.com/smpp/content~
                                        content=a772743846.

Badia:1998:PAC

[BV98] J. M. Badía and A. M. Vidal. Parallel algorithms to compute the eigenvalues and eigenvectors of symmetric Toeplitz matrices. Parallel Algorithms and Applications, 13(1):75–93, January 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776647677.

Becka:1999:BJSa

[BV99a] Martin Bečka and Marián Vajteršic. Block-Jacobi SVD algorithms for distributed memory systems. I. Hypercubes and rings. Parallel Algorithms and Applications, 13(3):265–287, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778232863.

Becka:1999:BJSb

[BV99b] Martin Bečka and Marián Vajteršic. Block-Jacobi SVD algorithms for distributed memory systems II: Meshes. Parallel Algorithms and Applications, 14(1):37–56, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777980764.

Baolin:1995:AMV

[BW95] Zhang Baolin and Li Wenzhi.

Age method with variable coefficients for parallel computing. Parallel Algorithms and Applications, 5(3–4):219–228, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400092.

Brown:1994:EPM

[BX94] Theodore Brown and Renbing Xiong. An EREW–PRAM multiple searching and merging algorithm. Parallel Algorithms and Applications, 3(1–2):83–88, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779202304.

Calinescu:1996:BSP

[Cal96] Radu Calinescu. Bulk synchronous parallel algorithms for conservative discrete event simulation. Parallel Algorithms and Applications, 9(1–2):15–38, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772744466.

Clematis:2002:PRA

[CC02] Andrea Clematis and Angelo Corana. Porting regular applications on heterogeneous workstation networks: Performance analysis and modeling. Parallel Algorithms and Applications, 17(3):205–226, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713601711.

Chen:1995:PCM

[CCY95] Yuh-Jiun Chen, Chin-Chen Chang, and Wei-Pang Yang. Parallel computation of the modular cascade exponentiation. Parallel Algorithms and Applications, 7(1–2):29–42, ???? 1995. CODEN PAAPEC.

ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743289.

Censor:1995:PSR

[CCZ95] Yair Censor, Emmanuel D. Chajakis, and Stavros A. Zenios. Parallelization strategies of a row-action method for multicommodity network flow problems. Parallel Algorithms and Applications, 6(2–3):179–205, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398625.

Criado:1997:AIM

[CD97] F. Criado and T. D. Davitashvili. On asynchronous iterative methods for the solution of non-linear equations and systems of equations for multiprocessors. Parallel Algorithms and Applications, 12(4):253–263, April 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776016225.

Chawla:1995:PSE

[CE95a] M. M. Chawla and D. J.

Evans. Parallel symmetric elimination for positive definite linear systems. Parallel Algorithms and Applications, 6(1):67–78, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397976.

Condurache:1995:IBC

[CE95b] Catalin Condurache and David J.

Evans. Implementation of a backpropagation-cauchy neural network training method on a parallel computer system. Parallel Algorithms and Applications, 5(1–2):69–77, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743842.

Calinescu:1996:QMA

[CE96] A. Calinescu and David J. Evans. A queueing modelling approach to clustered heterogeneous discrete event dynamic systems. Parallel Algorithms and Applications, 8(2):115–139, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314435.

Calinescu:1997:BSP

[CE97] Radu Calinescu and David J. Evans. Bulk-synchronous parallel algorithms for QR and QZ matrix factorisation. Parallel Algorithms and Applications, 11(1–2):97–112, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742764.

Celenk:1993:SAH

[Cel93] Mehmet Celenk. Speedup analysis of hypercube array processor for machine vision applications. Parallel Algorithms and Applications, 1(3):221–242, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398094.

Cappello:2000:PTO

[CES00] Peter Cappello, Ömer Eğecioğlu, and Chris Scheiman. Processortime-optimal systolic arrays. Parallel Algorithms and Applications, 15(3–4):167–199, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776716067.

Calamoneri:2002:MCC

[CFMP02] Tiziana Calamoneri, Irene Finocchi, Yannis Manoussakis, and Rossella Petreschi. On max cut in cubic graphs. Parallel Algorithms and Applications, 17(3):165–183, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713601712.

Chaudhuri:1996:FCM

[Cha96] Pranay Chaudhuri. Finding centers and medians of graphs in parallel. Parallel Algorithms and Applications, 9(1–2):111–118, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772744471.

Chaudhuri:1998:PIA

[Cha98] Pranay Chaudhuri. Parallel incremental algorithms for analyzing activity networks. Parallel Algorithms and Applications, 13(2):153–165, February 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776661631.

Chen:1997:OBS

[Che97] Lin Chen. Optimal bucket sorting and overlap representations. Parallel Algorithms and Applications, 10(3–4):249–269, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398798.

Christie:1995:CCC

[Chr95] Ruth Christie. The computational cost of clipping in parallel image processing. Parallel Algorithms and Applications, 5(1–2):149–160, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743847.

Chung:1994:SMC

[Chu94] Kuo-Liang L. Chung. Sorting on mesh-connected computers with segmented multiple buses. Parallel Algorithms and Applications, 4(1–2):71–75, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745100.

Choi:1993:OPA

[CJK93] Jung-Ju Choi, Chang-Sung Jeong, and Myung-Soo Kim. An optimal parallel algorithm for finding the smallest enclosing triangle on a mesh-connected computer. Parallel Algorithms and Applications, 1(2):115–126, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401501.

Cheng:2004:PSA

[CJP04] Jing-Ru C. Cheng, Mark T.

Jones, and Paul E. Plassmann. A portable software architecture for mesh-independent particle tracking algorithms. Parallel Algorithms and Applications, 19(2–3):145–161, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a714592652.

Chung:1995:GMD

[CLL95] Kuo-Liang Chung, R. C. T. Lee, and Yu-Chih Lin. A generalized mapping of 2-D array processor to linear array processor and its applications. Parallel Algorithms and Applications, 6(4):241–257, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772403912.

Carver:1996:PCH

[CLZ96] Doris L. Carver, Jigang Liu, and S. Q. Zheng. Planar convex hull algorithms on linear arrays. Parallel Algorithms and Applications, 10(1–2):59–70, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776220896.

Cruz:2003:PAM

[CM03] F. R. B. Cruz and G. R.

Mateus. Parallel algorithms for a multi-level network optimization problem. Parallel Algorithms and Applications, 18(3):121–137, September 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713812564.

Chopard:2000:PDE

[CPT00] Bastien Chopard, Olivier Pictet, and Marco Tomassini. Parallel and distributed evolutionary computation for financial applications. Parallel Algorithms and Applications, 15(1–2):15–36, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776636255.

Interface of parallel computing and statistics, economics, finance and optimization.

Chang:1993:BBD

[CS93] Mark Yan-Ming Chang and Martin H. Schultz. Bounds on block diagonal preconditioning. Parallel Algorithms and Applications, 1(2):141–164, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401503.

Chandrasekar:1994:SSD

[CS94] Srinivasan Chandrasekar and Pradip K. Srimani. A selfstabilizing distributed algorithm for all-pairs shortest path problem. Parallel Algorithms and Applications, 4(1–2):125–137, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745104.

Chen:1995:RSA

[CS95] Yuh-Shyan Chen and Jang-Ping Sheu. A reliable sorting algorithm on hypercube multicomputers. Parallel Algorithms and Applications, 5(3–4):165–186, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400089.

Chang:1996:TMR

[CT96] Pen-Yuang Chang and Jong-Chuang Tsay. Timespace mapping for regular arrays. Parallel Algorithms and Applications, 10(1–2):37–48, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776220894.

Celenk:1995:DCL

[CW95] Mehmet Celenk and Yang Wang.

Distributed computation in local area networks of workstations. Parallel Algorithms and Applications, 5(1–2):79–106, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743843.

Datta:2001:EPR

[Dat01] Amitava Datta. Efficient parallel range searching and partitioning algorithms. Parallel Algorithms and Applications, 16(4):301–317, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777327100.

Datta:2001:SSD

[DbT01] Ajoy K. Datta and Sébastien bastien Tixeuil. Self-stabilizing distributed sorting in tree networks. Parallel Algorithms and Applications, 16(1):1–15, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777308744.

Deng:2000:ROL

[DD00] Xiaotie Deng and Patrick Dymond. Randomized optimal list ranking on coarse-grained parallel computers with O(log p) communication phases. Parallel Algorithms and Applications, 14(3):165–173, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778235464.

Deep:1995:RSG

[DE95] K. Deep and D. J. Evans. The random search global optimization method for parallel computers. Parallel Algorithms and Applications, 5(3–4):269–282, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400096.

Douglas:1997:MSF

[DES97] Craig C. Douglas, Alexandre Ern, and Mitchell D. Smooke. Multigrid solution of flame sheet problems on serial and parallel computers. Parallel Algorithms and Applications, 10 (3–4):225–236, ???? 1997.

CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398785.

Dey:1994:LSN

[Dey94] S. K. Dey. Large-scale nonlinear parallel computations by perturbed functional iterations. Parallel Algorithms and Applications, 3(3–4):211–226, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167057.

Das:1994:EPA

[DFZ94] Sajal K. Das, Paul S. Fisher, and Hua Zhang. Efficient parallel algorithms for pattern recognition. Parallel Algorithms and Applications, 2(1–2):81–98, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745443.

Das:1994:EAM

[DHM94] Sajal K. Das, Wen-Bing Horng, and Gyo S. Moon. An efficient algorithm for managing a parallel heap. Parallel Algorithms and Applications, 4(3–4):281–299, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314738.

Djemame:2001:DSH

[Dje01] K. Djemame. Distributed simulation of high-level algebraic petri nets with limited capacity places. Parallel Algorithms and Applications, 16(3):207–241, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777320415.

Dogrusoz:1996:EAC

[DK96] U. Dogrusoz and M. S. Krishnamoorthy. Enumerating all cycles of a planar graph. Parallel Algorithms and Applications, 10(1–2):21–36, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776220893.

Dimov:1996:MCA

[DKKS96] I. Dimov, A. Karaivanova, H. Kuchen, and H. Stoltze. Monte Carlo algorithms for elliptic differential equations — a data parallel functional approach. Parallel Algorithms and Applications, 9(1–2):39–65, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772744467.

Diekmann:1996:CHS

[DLMS96] Ralf Diekmann, Reinhard Lüling, Burkhard Monien, and Carsten Spräner. Combining helpful sets and parallel simulated annealing for the graph-partitioning problem. Parallel Algorithms and Applications, 8(1):61–84, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a783554361.

Deshpande:1995:RAT

[DMSD95] A. Deshpande, S. Malhotra, M. H. Schultz, and C. C. Douglas. A rigorous analysis of time domain parallelism. Parallel Algorithms and Applications, 6(1):53–62, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397974.

Also available as Yale Computer Science Technical Report TR/DCS/996.

Denny:1997:IFD

[DOS97] W. M. Denny II, S. Olariu, and J. L. Schwing. The implementation of a fractal dimension estimation for texture discrimination on the dap. Parallel Algorithms and Applications, 10(3–4):211–224, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398784.

Douglas:1996:CMA

[Dou96] Craig C. Douglas. Caching in with multigrid algorithms: Problems in two dimensions. Parallel Algorithms and Applications, 9(3–4):195–204, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778707978.

Dikaiakos:1995:CSH

[DRS95] Marios D. Dikaiakos, Anne Rogers, and Ken Steiglitz. A comparison study of heuristics for mapping parallel algorithms to message-passing multiprocessors. Parallel Algorithms and Applications, 7(3–4):283–302, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772402452.

Dumitrescu:1994:FMM

[DRT94] B. Dumitrescu, J.-L. Roch, and D. Trystram. Fast matrix multiplications algorithms on MIMD architectures. Parallel Algorithms and Applications, 4(1–2):53–70, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745099.

Datta:1995:PAM

[DS95] Alak K. Datta and Ranjan K. Sen. A parallel algorithm for maximal matching based on depth first search. Parallel Algorithms and Applications, 5(3–4):161–164, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400020.

Datta:2000:IPA

[DS00] Alak Kumar Datta and Ranjan Kumar Sen. Improved parallel algorithm for maximal matching based on depth-first-search. Parallel Algorithms and Applications, 14(4):321–327, April 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778714136.

Datta:2004:TPM

[DS04] Alak Kumar Datta and Ranjan Kumar Sen. O(log_{4} n) time parallel maximal matching algorithm using linear number of processors. Parallel Algorithms and Applications, 19(1):19–32, March 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713646637.

Dehne:1995:APA

[DSVF95] Frank Dehne, Jörg-Rüdiger Sack, Natana Valiveti, and Boris Flach. Analog parallel algorithms for computational geometry. Parallel Algorithms and Applications, 5(1–2):1–14, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743838.

Darte:1997:OAK

[DV97] Alain Darte and Frédéric Vivien.

On the optimality of Allen and Kennedy’s algorithm for parallelism extraction in nested loops. Parallel Algorithms and Applications, 12(1–3):83–112, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a775978205.

Duan:1996:DAD

[DZ96] Zhaohui Duan and Zhaoqing Zhang. DPART: An automatic data partitioning system for distributed memory parallel machines. Parallel Algorithms and Applications, 9(3–4):205–212, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778707979.

Deren:1995:MRM

[DZE95] Wang Deren, Bai Zhongzhi, and D. J. Evans. Multisplitting relaxed methods for a class of system of nonlinear equations. Parallel Algorithms and Applications, 7(1–2):1–15, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743287.

Doumpos:2000:MSM

[DZP00] Michael Doumpos, Constantin Zopounidis, and Panos M. Pardalos. Multicriteria sorting methodology: Application to financial decision problems. Parallel Algorithms and Applications, 15(1–2):113–129, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776636259.

Interface of parallel computing and statistics, economics, finance and optimization.

Evans:1994:PIE

[EA94] D. J. Evans and R. Abdullah. The parallel implicit elimination (PIE) method for the solution of linear systems. Parallel Algorithms and Applications, 4(1–2):153–162, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745106.

Evans:1995:CQQ

[EA95] D. J. Evans and R. Abdullah. A comparison of the QR and QZ matrix factorisation methods on parallel computers. Parallel Algorithms and Applications, 7(1–2):43–52, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743290.

Evans:1997:SAS

[EA97] D. J. Evans and S. A. Amin.

Systolic algorithms for the solution of Toeplitz matrices. Parallel Algorithms and Applications, 12(4):321–339, April 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776016230.

Evans:1994:PRI

[EB94a] D. J. Evans and N. M. Bahoshy. Parallel Romberg integration of multidimensional integrals. Parallel Algorithms and Applications, 3(3–4):271–286, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167062.

Evans:1994:PSD

[EB94b] D. J. Evans and N. M. Bahoshy. The parallel solution of diagonal block bordered linear systems. Parallel Algorithms and Applications, 3(1–2):73–82, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779202303.

Evans:1995:UMN

[EB95] D. J. Evans and N. M. Bahoshy. A universal mapping for nested loops in parallel programming. Parallel Algorithms and Applications, 7(3–4):303–311, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772402453.

Evans:1994:LSA

[EC94] D. J. Evans and D. Caf. Linear systolic arrays for Toeplitz linear systems of equations. Parallel Algorithms and Applications, 2(3):155–172, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401562.

Evans:1995:DLB

[EC95] D. J. Evans and A. Calinescu. Dynamic load balancing algorithms for heavily and lightly loaded homogeneous distributed systems. Parallel Algorithms and Applications, 6(2–3):167–177, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398624.

Evans:1997:MIP

[EG97] David J. Evans and Marjan Gus̆ev. The magic of interlocking property: Fast systolic design. Parallel Algorithms and Applications, 10(3–4):195–209, ???? 1997. CODEN PAAPEC.

ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398783.

Evans:1994:CRA

[EJ94] D. J. Evans and Pragya Jain.

The coupled reduced alternating group explicit (CRAGE) method. Parallel Algorithms and Applications, 2(3):193–208, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401608.

Evangelinos:1996:PCB

[EK96] Constantinos Evangelinos and George Em Karniadakis. Parallel CFD benchmarks on Cray computers. Parallel Algorithms and Applications, 9 (3–4):273–298, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778707985.

Eshaghian:1993:HAI

[ENSS93] Mary M. Eshaghian, J. Greg Nash, Muhammad E. Shaaban, and David B. Shu. Heterogeneous algorithms for image understanding architecture. Parallel Algorithms and Applications, 1(4):273–284, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397699.

Evans:1997:PSS

[EO97] D. J. Evans and G. Okša.

Parallel solution of symmetric positive definite Toeplitz systems. Parallel Algorithms and Applications, 12(4):297–303, April 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776016228.

El-Rewini:1994:CCS

[ERA94] Hesham El-Rewini and Hesham H. Ali. On considering communication in scheduling task graphs on parallel processors. Parallel Algorithms and Applications, 3(3–4):177–191, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779167055.

El-Rewini:1995:OLU

[ERA95] Hesham El-Rewini and Hesham H. Ali. On optimal loop unrolling in two-processor scheduling. Parallel Algorithms and Applications, 7(1–2):107–119, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743294.

Egecioglu:1993:OPP

[ES93] Ömer Egecioğlu and Ashok Srinivasan. Optimal parallel prefix on mesh architectures. Parallel Algorithms and Applications, 1(3):191–209, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398092.

Evans:1994:BPP

[ES94] D. J. Evans and G. Spaletta. Block partitioning in the parallel recursive decoupling method. Parallel Algorithms and Applications, 3(3–4):331–347, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167075.

Evans:1996:PSC

[ES96] D. J. Evans and M. N. Sulaiman. Parallel simulation of character recognition problems using NEUCOMP2. Parallel Algorithms and Applications, 10(1–2):49–58, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776220895.

Entacher:2003:DPM

[ESSU03] Karl Entacher, Thomas Schell, Wolfgang Ch. Schmid, and Andreas Uhl. Defects in parallel Monte Carlo and quasi-Monte Carlo integration using the leap-frog technique. Parallel Algorithms and Applications, 18(1–2):13–26, March/ June 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713812543.

Algorithms and applications of parallel numerics (Bratislava, 2000).

Evans:1995:PDN

[ET95] D. J. Evans and L. P. Tay.

Parallel distributed neural networks for classification. Parallel Algorithms and Applications, 5(3–4):293–305, ???? 1995. CODEN PAAPEC.

ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772400098.

Erbas:1994:PMA

[ETN94] Cengiz Erbas, Murat M. Tanik, and V. S. S. Nair. Parallel memory allocation and data alignment in SIMD machines. Parallel Algorithms and Applications, 4(1–2):139–151, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745105.

Evans:1993:E

[Eva93] D. J. Evans. Editorial. Parallel Algorithms and Applications, 1(1):1, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772399752.

Evans:1997:PST

[Eva97] D. J. Evans. The parallel solution of tridiagonal systems by recursive striding. Parallel Algorithms and Applications, 10(3–4):161–164, January 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398780.

Evans:1997:LSA

[EW97] D. J. Evans and C. R. Wan. A linear systolic array for preconditioned conjugate gradient method. Parallel Algorithms and Applications, 11(3–4):227–247, ???? 1997. CODEN PAAPEC.

ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742292.

Evans:1994:QOD

[EY94a] D. J. Evans and P. Yalamov. The QZ orthogonal decomposition method. Parallel Algorithms and Applications, 2(4):263–276, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397560.

Evans:1994:QAC

[EY94b] D. J. Evans and W. S. Yousif.

The QZ algorithm for the calculation of the eigenvalues of a real matrix. Parallel Algorithms and Applications, 4(3–4):183–192, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314730.

Evans:1994:STL

[EY94c] D. J. Evans and W. S. Yousif.

Solving tridiagonal linear systems by the enhanced parallel stride of three reduction method. Parallel Algorithms and Applications, 2(3):149–154, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401561.

Ford:2003:FPF

[FCF03] Judith M. Ford, Ke Chen, and Neville J. Ford. Flexible parallelization of fast wavelet transforms. Parallel Algorithms and Applications, 18 (4):155–169, December 2003.

CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713601714.

Ferreira:1996:GBI

[FGS96] A. Ferreira, A. Goldman vel Lejbman, and S. W. Song. Gossiping in bus interconnection networks. Parallel Algorithms and Applications, 8 (3–4):309–331, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776632238.

Ferrante:1997:CTA

[FGS97a] Jeanne Ferrante, Dirk Grunwald, and Harini Srinivasan. Compile-time analysis and optimization of explicitly parallel programs. Parallel Algorithms and Applications, 12 (1–3):21–56, October 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a775978203.

Folwell:1997:PAI

[FGS97b] Nathan Folwell, Sumanta Guha, and Ichiro Suzuki. A practical algorithm for integer sorting on a mesh-connected computer. Parallel Algorithms and Applications, 12(4):265–278, April 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776016226.

Ferreira:19xx:BBI

[FGSxx] A. Ferreira, A. Goldman vel Lejbman, and S. W. Song. Broadcasting in bus interconnection networks. Parallel Algorithms and Applications, ?? (??):??, ???? 19xx. CODEN PAAPEC. ISSN 1063-7192. See also the author’s paper in Parallel Processing: CONPAR 94 - VAPP VI. Third Joint International Conference on Vector and Parallel Processing Proceedings, Springer-Verlag, p. xvi + 889, pp. 797–807 (1994).

Fidanova:2001:LAC

[Fid01] Stefka Fidanova. Linear array for a class of non uniform recurrence equations. Parallel Algorithms and Applications, 16(4):273–281, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777327098.

Fouliras:1998:PDD

[FM98] P. E. Fouliras and K. G. Margaritis. A parallel demanddriven simulation algorithm for SIMD computers. Parallel Algorithms and Applications, 13(1):53–73, January 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776647676.

Francomano:1995:RFV

[FML95] E. Francomano, A. Tortorici Macaluso, and C. Lodato. A recurrence-free variant of Strassen’s algorithm on hypercube. Parallel Algorithms and Applications, 5(3–4):241–249, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772400094.

Fulton:1994:PCS

[FPS94] Charles Fulton, Steven Pruess, and William Shoaff. Parallel computation of Sturm– Liouville spectral density functions. Parallel Algorithms and Applications, 4(1–2):41–51, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745098.

Friedl:1996:RRB

[FT96] Katalin Friedl and Shi-Chun Tsai. Recycling random bits in parallel. Parallel Algorithms and Applications, 8(1):85–94, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772400361.

Fagg:1995:IPP

[FW95] Graham E. Fagg and Shirley A. Williams. Improved program performance using a cluster of workstations. Parallel Algorithms and Applications, 6 (2–3):233–236, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398643.

Gravvanis:2004:JPA

[GA04] George A. Gravvanis and Hamid R. Arabnia. The Journal of Parallel Algorithms and Applications: Special issue on parallel and distributed algorithms. Parallel Algorithms and Applications, 19(2–3):77–78, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a714592648.

Gengler:1994:PBF

[GC94] M. Gengler and G. Coray. A parallel best-first B&B algorithm and its axiomatization. Parallel Algorithms and Applications, 2(1–2):61–80, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745442.

Gusev:1993:FMV

[GE93] Marjan Gus̆ev and David J.

Evans. The fastest matrix vector multiplication. Parallel Algorithms and Applications, 1(1):57–67, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772399780.

Gusev:1994:FTS

[GE94] M. Gus̆ev and D. J. Evans.

Folding transformations on systolic and VLSI processor arrays. Parallel Algorithms and Applications, 4(3–4):239–274, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777314736.

Gusev:1996:IPL

[GE96] Marjan Gus̆ev and David J. Evans. Interlocking properties of the linear data dependence method. Parallel Algorithms and Applications, 8(2):97–114, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777314434.

Gerbessiotis:2003:TTB

[Ger03] Alexandros V. Gerbessiotis. Trinomial-tree based parallel option price valuations. Parallel Algorithms and Applications, 18(4):181–196, December 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713646632.

Gravvanis:2004:DNE

[GGM04] G. A. Gravvanis, K. M. Giannoutakis, and N. M. Missirlis. A distributed normalized explicit preconditioned conjugate gradient method. Parallel Algorithms and Applications, 19(2–3):163–174, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a714592653.

Galley:1996:PAS

[GI96] Clive N. Galley and Costas S. Iliopoulos. A PRAM algorithm for a special case of the set partition problem. Parallel Algorithms and Applications, 10(1–2):105–109, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776220901.

Gasparo:1995:PMQ

[GMP95] M. G. Gasparo, B. Morini, and A. Papini. A parallel method for quasilinear problems. Parallel Algorithms and Applications, 6(1):79–86, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398008.

Golik:2000:PSP

[Gol00]  Wojciech L. Golik.  Paral-

lel solvers for planar elliptic grid generation equations. Parallel Algorithms and Applications, 14(3):175–186, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778235465.

Guarracino:1994:PBL

[GP94] Mario Rosario Guarracino and Francesca Perla. A parallel block Lanczos algorithm for distributed memory architectures. Parallel Algorithms and Applications, 4(3–4):211–221, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314732.

Galligani:1994:PPB

[GR94] E. Galligani and V. Ruggiero.

A polynomial preconditioner for block tridiagonal matrices. Parallel Algorithms and Applications, 3(3–4):227–237, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167058.

Garg:1994:OMM

[GSA94] Sharad Garg, Howard A. Sholl, and Reda A. Ammar. Optimizing the matrix multiplication parallel algorithms on a distributed-memory MIMD multiprocessor. Parallel Algorithms and Applications, 2(4):291–303, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397562.

Guivarch:1997:SAP

[GSBM97] R. Guivarch, P. Spiteri, H. C. Boisson, and J. C. Miellou. Schwarz alternating parallel algorithm applied to incompressible flow computation in vorticity stream function formulation. Parallel Algorithms and Applications, 11(3–4):205–225, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742291.

Guyennet:1995:MSD

[GST95] Hervé Guyennet, François Spies, and Michel Tréhel. Modeling and simulation of dynamic load balancing using queueing theory. Parallel Algorithms and Applications, 5(3–4):199–218, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400091.

Garey:2002:PDC

[GSZ02] L. E. Garey, R. E. Shaw, and J. Zhang. A parallel divide and conquer algorithm for non symmetric tridiagonal Toeplitz systems using conjugate gradient. Parallel Algorithms and Applications, 17(4):249–262, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713941499.

Gumel:1997:PSM

[GTA97] A. B. Gumel, E. H. Twizell, and M. A. Arigu. Parallel L_{0}- stable methods for the multidimensional diffusion equation. Parallel Algorithms and Applications, 11(1–2):13–25, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742759.

Guha:1998:OMA

[Guh98]  Sumanta Guha.  Optimal

mesh algorithms for proximity and visibility problems in simple polygons. Parallel Algorithms and Applications, 13 (2):167–185, February 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776661632.

Hartley:1997:PSE

[HB97] Joanna K. Hartley and Andrzej Bargiela. Parallel state estimation with confidence limit analysis. Parallel Algorithms and Applications, 11(1–2):155–167, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742767.

Hassen:1997:HRS

[HBT97] Saniya Ben Hassen, Henri E. Bal, and Andrew S. Tanenbaum. Hawk: a runtime system for partitioned objects. Parallel Algorithms and Applications, 12(1–3):205–230, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a775978210.

Hayashi:2002:SPP

[HC02] Yasushi Hayashi and Murray Cole. Static performance prediction of skeletal parallel programs. Parallel Algorithms and Applications, 17(1):59–84, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777803178.

Hsu:1993:OPP

[HCL93] F. R. Hsu, R. C. Chang, and R. C. T. Lee. Optimal parallel preprocessing algorithms for testing weak visibility of polygons from segments. Parallel Algorithms and Applications, 1(2):83–98, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401485.

He:1995:SDD

[HE95] Qiming He and David J. Evans.

Symmetric domain decomposition method for large scale and nonlinear evolution systems. Parallel Algorithms and Applications, 7(3–4):161–175, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772402404.

Han:1996:SAN

[HE96a] Yongfei Han and David J.

Evans. On systolic algorithms and neural nets for inference. Parallel Algorithms and Applications, 8(2):169–175, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314442.

He:1996:MSP

[HE96b] Qiming He and D. J. Evans.

Monotone-Schwarz parallel algorithm for nonlinear elliptic equations. Parallel Algorithms and Applications, 9(3–4):173–184, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778707976.

Henrich:1996:LML

[Hen96] Dominik Henrich. The liquid model load balancing method. Parallel Algorithms and Applications, 8(3–4):285–307, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776632237.

Special Issue on Algorithms for Enhanced Mesh Architectures.

He:1995:CGE

[HKC95] Jun He, Li-Shan Kang, and Yong-Jun Chen. Convergence of genetic evolution algorithms for optimization. Parallel Algorithms and Applications, 5(1–2):37–56, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743840.

Hafsteinnson:1994:SLS

[HLM94] Hjálmtyýr Hafsteinnson, Roni Levkovitz, and Gautam Mitra. Solving large scale linear programming problems using an interior point method on a massively parallel SIMD computer. Parallel Algorithms and Applications, 4(3–4):301–316, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777314739.

Hu:1995:DIP

[HLN95] Jie Hu, Lei Li, and Tadao Nakamura. A divide-and-inner product parallel algorithm for polynomial evaluation. Parallel Algorithms and Applications, 6(1):63–66, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772397975.

Hu:1997:AMN

[HNL97a] Jie Hu, Tadao Nakamura, and Lei Li. Asynchronous monotone Newton iterative method on distributed computers. Parallel Algorithms and Applications, 12(4):341–348, April 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776016231.

Hu:1997:CCS

[HNL97b] Jie Hu, Tadao Nakamura, and Lei Li. Convergence, complexity and simulation of monotone asynchronous iterative method for computing fixed point[s] on a distributed computer. Parallel Algorithms and Applications, 11(1–2):1–11, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742757.

Holmgren:2002:PPS

[HNRW02] Sverker Holmgren, Markus Nordén, Jarmo Rantakokko, and Dan Wallin. Performance of PDE solvers on a self-optimizing NUMA architecture. Parallel Algorithms and Applications, 17(4):285–299, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713941501.

Horng:1994:PDC

[Hor94] S. J. Horng. Parallel dominator computation on a RAP. Parallel Algorithms and Applications, 3(1–2):1–8, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779202298.

Heydemann:1996:ECB

[HSO96] M.-C. Heydemann, D. Sotteau, and J. Opatrny. Embeddings of complete binary trees into extended grids with edge-congestion 1∗. Parallel Algorithms and Applications, 8(3–4):333–354, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776632239.

Haridas:2004:FIC

[HZ04] Satchidanand G. Haridas and Sotirios G. Ziavras. FPGA implementation of a Cholesky algorithm for a shared-memory multiprocessor architecture. Parallel Algorithms and Applications, 19(4):211–226, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a714592657.

Jianxin:1996:LDP

[JDMW96] Xiong Jianxin, Wang Dingxing, Shen Meiming, and Zheng Weimin. On-line debugging of parallel programs. Parallel Algorithms and Applications, 9(3–4):227–236, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a778707981.

Johnson:1994:PAS

[JLT94] Theodore Johnson, Panos E. Livadas, and Sunjay E. Talele. A parallel algorithm for surface triangulation. Parallel Algorithms and Applications, 3(1–2):9–30, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779202299.

Jenq:1993:HRM

[JS93] Jing-Fu Jenq and Sartaj Sahni.

Histogramming on a reconfigurable mesh computer. Parallel Algorithms and Applications, 1(3):179–190, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398091.

Kontoghiorghes:1993:SUD

[KC93] E. J. Kontoghiorghes and M. R. B. Clarke. Solving the updated and downdated ordinary linear model on massively parallel SIMD systems. Parallel Algorithms and Applications, 1(3):243–252, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398095.

Kang:1996:APA

[KCMC96] Lishan Kang, Yuping Chen, Iain MacLeod, and Lujuan Chen. Asynchronous parallel algorithms based on DDM on CM-5 (I) for solving linear elliptic PDE’s. Parallel Algorithms and Applications, 9(3–4):161–172, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778707975.

Koc:1994:ESL

[KGB94]  Ç. K. Koç, A. Güvenç, and

B. Bakkaloğlu. Exact solution of linear equations on distributed-memory multiprocessors. Parallel Algorithms and Applications, 3(1–2):135–143, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779202347.

Kharab:2001:ESM

[Kha01]  A. Kharab.  Eigenpairs of

symmetric matrices using the quadratic method and the method of subdefinite calculations. Parallel Algorithms and Applications, 16(4):319–326, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777327101.

Kiper:1996:PPE

[Kip96] Ayse Kiper. Parallel polynomial evaluation by decoupling algorithm. Parallel Algorithms and Applications, 9(1–2):145–152, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772744476.

Kim:1995:OPMa

[KJ95a] Myung-Ho Kim and Chang-Sung Jeong. An optimal parallel matching algorithm for a convex bipartite graph on a mesh-connected computer. Parallel Algorithms and Applications, 5(1–2):15–35, ???? 1995. CODEN PAAPEC.

ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743839.

Kim:1995:OPMb

[KJ95b] Myung-Ho Kim and Chang-Sung Jeong. An optimal parallel matching algorithm for a convex bipartite graph on a mesh-connected computer. Parallel Algorithms and Applications, 6(2–3):129–141, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398622.

Kaskalis:1996:RST

[KM96] T. H. Kaskalis and K. G. Margaritis. A reconfigurable systolic torus for iterative matrix operations. Parallel Algorithms and Applications, 10(1–2):79–91, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776220899.

Kim:1996:PSS

[KMBK96] Taegeun Kim, Kyeongdeok Moon, Nanjoo Ban, and Jungkwon Kim. PPTran: Source to source translator for High Performance Fortran. Parallel Algorithms and Applications, 9(3–4):213–225, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778707980.

Kontoghiorghes:1995:NPS

[Kon95] Erricos J. Kontoghiorghes. New parallel strategies for block updating the QR decomposition. Parallel Algorithms and Applications, 5(3–4):229–239, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400093.

Kontoghiorghes:2000:PGS

[Kon00] Erricos John Kontoghiorghes. Parallel Givens sequences for solving the general linear model on a EREW PRAM. Parallel Algorithms and Applications, 15(1–2):57–75, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776636257.

Interface of parallel computing and statistics, economics, finance and optimization.

Kadalbajoo:1995:PMS

[KR95] Mohan K. Kadalbajoo and A. Appaji Rao. Parallel method for solving singularly perturbed boundary value problems. Parallel Algorithms and Applications, 7(1–2):121–131, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743295.

Kadalbajoo:1996:PAA

[KR96] M. K. Kadalbajoo and A. Appaji Rao. Parallel age algorithm for parabolic equations with a small parameter. Parallel Algorithms and Applications, 8(2):155–168, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314437.

Kadalbajoo:1997:PAA

[KR97] Mohan K. Kadalbajoo and A. Appaji Rao. Performance analysis of age algorithm for singular perturbation problems. Parallel Algorithms and Applications, 10(3–4):315–318, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398802.

Konig:1998:AGA

[KRT98] J.-C. König, P. S. Rao, and D. Trystram. Analysis of gossiping algorithms with restricted buffers. Parallel Algorithms and Applications, 13 (2):117–133, February 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776661629.

Kuzyurin:1995:PAF

[KS95] Nikolai N. Kuzyurin and Leonid V.

Shabanov. A parallel algorithm for fixed-dimensional linear programming. Parallel Algorithms and Applications, 6(1):17–24, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397971.

Kim:1997:OPM

[KS97] Hiryoung Kim and Alan P.

Sprague. Optimal parallel matching on bipartite permutation graphs. Parallel Algorithms and Applications, 10 (3–4):165–176, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398781.

Langston:1996:PAI

[Lan96] Michael A. Langston. Parallel algorithms, an introduction to the special issue. Parallel Algorithms and Applications, 8(1):1–2, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772400252.

Lau:2002:LDP

[Lau02] Francis C. M. Lau. The load distribution problem in a processor ring. Parallel Algorithms and Applications, 17(3):185–204, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713601710.

Lai:1997:SPP

[LC97] Guan-Joi Lai and Cheng Chen. Scheduling parallel program tasks with non-negligible intertask communications on to NUMA multiprocessor systems. Parallel Algorithms and Applications, 12(1–3):165–184, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a775978208.

Lee:2004:HJP

[LCFL04] Han-Ku Lee, Bryan Carpenter, Geoffrey Fox, and Sang Boem Lim. HP Java: Programming support for high-performance Grid-enabled applications. Parallel Algorithms and Applications, 19(2–3):175–193, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a714592654.

Li:1995:PIS

[LE95] Changjun Li and D. J. Evans.

Parallel implementation for the symmetric eigenvalue problems on a Meiko i860 computer. Parallel Algorithms and Applications, 6(2–3):229–232, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398642.

Li:1997:STS

[Li97] Keqin Li. Solving the traveling salesman problem using efficient randomized parallel approximation algorithms. Parallel Algorithms and Applications, 10(3–4):271–281, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398799.

Li:2004:FSP

[Li04] Keqin Li. Fast and scalable parallel matrix computations with reconfigurable pipelined optical buses. Parallel Algorithms and Applications, 19(4):195–209, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a714592656.

Levene:1993:MLF

[LL93] Mark Levene and George Loizou. A modal logic formalism for distributed and parallel knowledge bases. Parallel Algorithms and Applications, 1(1):11–27, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772399754.

Luce:1998:ASA

[LM98] Guillaume Luce and Jean Frédéric Myoupo. Application-specific array processors for the longest common subsequence problem of three sequences. Parallel Algorithms and Applications, 13(1):27–52, January 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776647675.

Li:1996:MRM

[LMR96] Zhiyong Li, Peter H. Mills, and John H. Reif. Models and resource metrics for parallel and distributed computation. Parallel Algorithms and Applications, 8(1):35–59, December 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a783554362.

Li:1995:CAI

[LN95] Lei Li and Tadao Nakamura. The convergence of asynchronous iterations for the fixed point of a splitting operator. Parallel Algorithms and Applications, 7(3–4):229–235, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772402408.

Lin:1995:COE

[LO95] Rong Lin and Stephan Olariu.

A cost-optimal EREW breadthfirst algorithm for ordered trees, with applications. Parallel Algorithms and Applications, 5 (3–4):187–197, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400090.

Lin:1994:EEA

[LOSZ94] R. Lin, S. Olariu, J. L. Schwing, and J. Zhang. An efficient EREW algorithm for minimum path cover and Hamiltonicity on cographs. Parallel Algorithms and Applications, 2(1–2):99–113, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745444.

Liou:1997:PAT

[LPW97] Jing-Chiou Liou, Michael A.

Palis, and David S. L. Wei. Performance analysis of task clustering heuristics for scheduling static DAGs on multiprocessor system. Parallel Algorithms and Applications, 12(1–3):185–203, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a775978209.

Liszka:1996:PKT

[LQ96] K. Liszka and A. Quesada. On the parallel k th extension of the Sieve of Eratosthenes. Parallel Algorithms and Applications, 10(1–2):111–125, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776220902.

Lee:1997:OSA

[LTCT97] Wei-Ping Lee, Jong-Chuang Tsay, Hsien-Sheng Chen, and Tai-Jyi Tseng. An optimal systolic algorithm for the set partitioning problem. Parallel Algorithms and Applications, 10(3–4):301–313, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398801.

Lowekamp:1996:CAP

[LWC96] Bruce B. Lowekamp, Layne T. Watson, and Mark S. Cramer. The cellular automata paradigm for the parallel solution of heat transfer problems. Parallel Algorithms and Applications, 9(1–2):119–130, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772744472.

Lenders:2002:EBP

[LX02] Patrick Lenders and Jingling Xue. Eigenvectors-based parallelisation of nested loops with affine dependences. Parallel Algorithms and Applications, 17(3):227–248, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713601713.

Lin:2002:DFS

[LY02] Yen-Chun Lin and Jih-Wei Yeh.

Deriving a fast systolic algorithm for the longest common subsequence problem. Parallel Algorithms and Applications, 17(1):1–18, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777803175.

Megson:1994:SMI

[MAD94] G. M. Megson, V. N. Aleksandrov, and I. T. Dimov. Systolic matrix inversion using a Monte Carlo method. Parallel Algorithms and Applications, 3(3–4):311–330, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167064.

Margaritis:1994:ANN

[MAGE94] K. G. Margaritis, M. Adamopoulos, K. Goulianas, and D. J. Evans. Artificial neural networks and iterative linear algebra methods. Parallel Algorithms and Applications, 3(1–2):31–44, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779202300.

Mayr:1996:SIO

[May96]  Ernst W. Mayr.  Schedul-

ing interval orders in parallel. Parallel Algorithms and Applications, 8(1):21–34, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772400358.

Merry:1995:CTA

[MB95] Mark S. Merry and Johnnie Baker. A constant time algorithm for the channel assignment problem using the reconfigurable mesh. Parallel Algorithms and Applications, 6(4):259–271, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772403913.

Mohamed:2004:LCL

[MB04] Ahmed S. Mohamed and Veysel S. Baydogan. A localityconscious load-balancer based on negotiations in dynamic unstructured mesh computations. Parallel Algorithms and Applications, 19(2–3):97–125, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a714592650.

Makhaniok:1995:ECS

[MCMN95] M. Makhaniok, V. Cherniavsky, R. Männer, and K.-H. Noffz. Extremal codes for speed-up of distributed parallel arbitration. Parallel Algorithms and Applications, 6(1):1–16, ???? 1995. CODEN PAAPEC.

ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397970.

Murphy:1997:PCP

[MCSW97] K. Murphy, M. Clint, M. Szularz, and J. Weston. The parallel computation of partial eigensolutions using a modified Lanczos method. Parallel Algorithms and Applications, 11(3–4):299–323, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742296.

Margaritts:1994:SGS

[ME94a] K. G. Margaritts and D. J.

Evans. On systolic general system solution and general matrix inversion. Parallel Algorithms and Applications, 3 (3–4):193–210, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167056.

Martins:1994:PMG

[ME94b] M. M. Martins and D. J.

Evans. The parallel multisplitting generalized iterative methods. Parallel Algorithms and Applications, 2(4):251–262, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397559.

Margaritis:1995:FTS

[ME95] K. G. Margaritis and D. J.

Evans. Folding techniques for systolic iterations. Parallel Algorithms and Applications, 7(1–2):87–105, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743293.

Margaritis:1997:VPA

[ME97] K. G. Margaritis and D. J. Evans. A VLSI processor array for flexible string matching. Parallel Algorithms and Applications, 11(1–2):45–60, May 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772742761.

Mohanty:1999:BIM

[ME99] R. K. Mohanty and D. J. Evans. Block iterative methods for one dimensional nonlinear biharmonic problems on a parallel computer. Parallel Algorithms and Applications, 13(3):239–263, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a778232862.

Megson:1993:DUR

[Meg93] G. M. Megson. The derivation of Uniform Recurrence Equations for the Knapsack Problem. Parallel Algorithms and Applications, 1(2):127–140, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772401502.

Mittelmann:1996:PMC

[Mit96] H. D. Mittelmann. Parallel multisplittings for constrained optimization. Parallel Algorithms and Applications, 9(1–2):91–99, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772744469.

Mansour:1997:PSA

[MKD97] Nashat Mansour, Jalal Kawash, and Hassan Diab. Properties of simulated annealing and genetic algorithms for mapping data to multicomputers. Parallel Algorithms and Applications, 12(4):279–296, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776016227.

Murthy:1994:TSE

[MM94] K. N. Balasubramanya Murthy and C. Siva Ram Murthy. Twosided elimination algorithm for parallel solution of linear equations using Householder reductions. Parallel Algorithms and Applications, 3(3–4):287–309, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167063.

Murthy:1995:NMG

[MM95] K. N. Balasubramanya Murthy and C. Siva Ram Murthy. A new modified Gram–Schmidt orthogonal matrix factorization based algorithm for parallel solution of linear equations. Parallel Algorithms and Applications, 6(1):39–52, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772397973.

Manjunathaiah:2004:TRA

[MM04] M. Manjunathaiah and G. M. Megson. Tools for regularizing array designs. Parallel Algorithms and Applications, 19(1):51–75, March 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713646640.

Mathioudakis:1996:MPI

[MPS96] E. N. Mathioudakis, E. P. Papadopoulou, and Yiannis G. Saridakis. Mapping parallel iterative algorithms for PDE computations on a distributed memory computer. Parallel Algorithms and Applications, 8(2):141–154, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777314436.

Metallinos:1994:ENA

[MRS94] S. N. Metallinos, D. I. Reisis, and G. I. Stassinopoulos. An efficient network analyser based on linear array architecture. Parallel Algorithms and Applications, 2(1–2):139–147, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745447.

Matias:1996:FEM

[MS96] Yossi Matias and Assaf Schuster. Fast, efficient mutual and self simulations for shared memory and reconfigurable mesh. Parallel Algorithms and Applications, 8(3–4):195–221, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776632201.

Manolakos:2000:SSP

[MS00a] Elias S. Manolakos and Haris M.

Stellakis. Systematic synthesis of parallel architectures for the real-time estimation of higher order statistical moments. Parallel Algorithms and Applications, 15(1–2):77–111, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776636258.

Interface of parallel computing and statistics, economics, finance and optimization.

Myoupo:2000:EPA

[MS00b] Jean-Frédéric Myoupo and David Semé. Efficient parallel algorithms for the LIS and LCS problems on BSR model using constant number of selections. Parallel Algorithms and Applications, 14(3):187–202, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778235466.

Michaelson:2001:NAS

[MSBK01] Greg Michaelson, Norman Scaife, Paul Bristow, and Peter King. Nested algorithmic skeletons from higher order functions.

Parallel Algorithms and Applications, 16(3):181–206, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777320411.

Maniatty:19xx:SME

[MSCSxx] Bill Maniatty, Boleslaw K. Szymanski, Tom Caraco, and Balaram Sinharoy. Spatial modeling of epidemics on parallel machines. Parallel Algorithms and Applications, ??(??):??, ???? 19xx. CODEN PAAPEC. ISSN 1063-7192.

Moriarty:1994:HSM

[MST94] K. J. M. Moriarty, S. Sanielevici, and T. Trappenberg. High speed Monte Carlo simulations on vector parallel computers. Parallel Algorithms and Applications, 3(3–4):165–175, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779167014.

Missirlis:1995:PBM

[MT95] N. M. Missirlis and F. I. Tjaferis. Parallel block methods for solving linear systems. Parallel Algorithms and Applications, 5(1–2):57–68, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743841.

Missirlis:1997:HMS

[MT97] N. M. Missirlis and F. I. Tjaferis. The Huard method on a shared memory MIMD computer. Parallel Algorithms and Applications, 11(3–4):249–272, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742293.

Mocnik:2003:ILB

[MTR03] Jaka Moćnik, Roman Trobec, and Borut Robić. Integration of load balancing in a CORBA environment. Parallel Algorithms and Applications, 18(1–2):99–105, March/ June 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713812548.

McMahon:2000:DGA

[MW00] Matthew T. McMahon and Layne T. Watson. A distributed genetic algorithm with migration for the design of composite laminate structures. Parallel Algorithms and Applications, 14(4):329–362, April 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778714137.

Nakano:1996:CCH

[Nak96]  Koji Nakano.  Computation

of the convex hull for sorted points on a reconfigurable mesh. Parallel Algorithms and Applications, 8(3–4):243–250, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776632234.

Nasuto:1999:CAS

[NB99] S. Nasuto and M. Bishop. Convergence analysis of stochastic diffusion search. Parallel Algorithms and Applications, 14(2):89–107, February 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777957921.

Nikolopoulos:1994:FPA

[ND94] Stavros D. Nikolopoulos and Stylianos D. Danielopoulos. Fast parallel algorithms for finding cutpoints and bridges of undirected graphs. Parallel Algorithms and Applications, 3(1–2):45–55, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779202301.

Nikolopoulos:1996:PBF

[Nik96] Stavros D. Nikolopoulos. Parallel block-finding using distance matrices. Parallel Algorithms and Applications, 9(1–2):1–13, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772744465.

Osiakwan:1994:EPA

[OA94] Constantine N. K. Osiakwan and Selim G. Akl. An efficient parallel algorithm for the assignment problem on the plane. Parallel Algorithms and Applications, 4(3–4):193–210, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314731.

Osiakwan:1995:MWP

[OA95] Constantine N. K. Osiakwan and Selim G. Akl. The maximum weight perfect matching problem for complete weighted graphs is in PC*. Parallel Algorithms and Applications, 6(2–3):143–166, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398623.

Othman:2004:PFP

[OAE04] M. Othman, A. R. Abdullah, and D. J. Evans. A parallel four points modified explicit group algorithm on shared memory multiprocessors. Parallel Algorithms and Applications, 19(1):1–9, March 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713646636.

Oliveira:2000:ADP

[OBHS00] S. Oliveira, L. Borges, M. Holzrichter, and T. Soma. Analysis of different partitioning schemes for parallel Gram–Schmidt algorithms. Parallel Algorithms and Applications, 14(4):293–320, April 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778714135.

Olariu:1996:F

[Ola96]  Stephan Olariu.  Foreword.

Parallel Algorithms and Applications, 8(3–4):177–178, ???? 1996. CODEN PAAPEC.

ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776632199.

Olariu:1993:SSA

[OSS^{+}93] S. Olariu, J. L. Schwing, W. Shen, L. Wilson, and J. Zhang. A simple selection algorithm for reconfigurable meshes. Parallel Algorithms and Applications, 1(1):29–41, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772399755.

Oksa:2003:SBJ

[OV03] Gabriel Okša and Marián Vajteršic. A systolic block-Jacobi SVD solver for processor meshes. Parallel Algorithms and Applications, 18(1–2):49–70, March/ June 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a716776278.

Algorithms and applications of parallel numerics (Bratislava, 2000).

Pavel:1996:MOU

[PA96] Sandy Pavel and Selim G. Akl. Matrix operations using arrays with reconfigurable optical buses. Parallel Algorithms and Applications, 8(3–4):223–242, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776632233.

Papadopoulos:1994:NHA

[Pap94] Constantinos V. Papadopoulos. A new hashing algorithm for parallel processors. Parallel Algorithms and Applications, 4(3–4):223–237, November 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314733.

This is a plagiarized arti-

cle. See http://www.sics.se/
europar95/plagiarism.html

for details. The original work from which the material in this paper was stolen is due to Thomas J. Sheffler and Randal E. Bryant, CMU report MCU-CS-92-172.

Papadopoulos:1995:EMA

[Pap95] Constantinos V. Papadopoulos. Evaluation of multigrid algorithms on message-passing multiprocessors. Parallel Algorithms and Applications, ??(??): ??, September 1995. CODEN PAAPEC. ISSN 1063-7192. This is believed to be a plagiarized ar-

ticle. See http://www.sics.se/
europar95/plagiarism.html
and http://www-db.stanford.
edu/~shiva/SCAM/end.html for

details. The original work from which the material in this paper was stolen has not yet been identified.

Papay:1996:PPP

[PAZN96] J. Papay, T. J. Atherton, M. J.

Zemerly, and G. R. Nudd. Performance prediction of parallel self consistent field computation. Parallel Algorithms and Applications, 10(1–2):127–143, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776220903.

Pal:1995:PAD

[PB95] Madhumangal Pal and G. P. Bhattacharjee. The parallel algorithms for determining edgepacking and efficient edge dominating sets in interval graphs. Parallel Algorithms and Applications, 7(3–4):193–207, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772402406.

Pagliaro:2000:PCS

[PB00] A. Pagliaro and U. Becciani. A parallel code for structure detection. Parallel Algorithms and Applications, 14(??):1–20, July 2000. CODEN PAAPEC. ISSN 1063-7192.

Pean:2001:EPT

[PC01] Der-Lin Pean and Cheng Chen. Effective parallelization techniques for loop nests with nonuniform dependences. Parallel Algorithms and Applications, 16(1):37–64, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777308746.

Punyamurtula:1997:CTP

[PCJR97] Swamy Punyamurtula, Vipin Chaudhary, Jialin Ju, and Sumit Roy. Compile time partitioning of nested loop iteration spaces with non-uniform dependences. Parallel Algorithms and Applications, 12(1–3):113–141, January 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a775978206.

Petrini:2001:IRU

[PF01] Fabrizio Petrini and Wu-Chun Feng. Improved resource utilization with buffered coscheduling. Parallel Algorithms and Applications, 16(2–3):123–144, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777319806.

Piano:2003:PSL

[Pia03] Cristiana Piano. The problem of small and large matrices in parallel matrix multiplication. Parallel Algorithms and Applications, 18(1–2):71–89, March/ June 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713812546.

Algorithms and applications of parallel numerics (Bratislava, 2000).

Pini:1994:TFL

[Pin94] G. Pini. Transpose-free Lanczostype schemes on transputer network. Parallel Algorithms and Applications, 3(3–4):249–260, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167060.

Pineo:1997:ECS

[Pin97] Patricia Prather Pineo. The efficient creation of single assignment forms from unstructured code. Parallel Algorithms and Applications, 12 (1–3):143–163, February 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a775978207.

Pini:2002:LER

[Pin02] Giorgio Pini. Leftmost eigenvalue of real and complex sparse matrices on parallel computer using approximate inverse preconditioning. Parallel Algorithms and Applications, 17(1):41–58, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777803177.

Peng:1994:OLS

[PL94] Shietung Peng and Win-Tsung Lo. The optimal location of a structured facility in a tree network. Parallel Algorithms and Applications, 2(1–2):43–60, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745441.

Plaks:1999:EMR

[Pla99] Toomas P. Plaks. Efficient mapping reductions using isoplanes on the polytope model. Parallel Algorithms and Applications, 13(4):321–343, April 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776633663.

Plaks:2000:ARA

[Pla00] Toomas P. Plaks. Advanced regular array design. Parallel Algorithms and Applications, 15(3–4):131–135, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776716065.

Pan:1998:FNN

[PLZ98] Yi Pan, Keqin Li, and Si-Qing Zheng. Fast nearest neighbor algorithms on a linear array with a reconfigurable pipelined bus system. Parallel Algorithms and Applications, 13(1):1–25, January 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776647674.

Padmini:1999:LALa

[PMJ99a] M. V. Padmini, B. B. Madan, and B. N. Jain. A linear array for large sparse matrix operations — I. Cholesky factorization. Parallel Algorithms and Applications, 13(3):187–215, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778232860.

Padmini:1999:LALb

[PMJ99b] M. V. Padmini, B. B. Madan, and B. N. Jain. A linear array for large sparse matrix operations — II. triangular system solvers and matrix multiplication. Parallel Algorithms and Applications, 13(3):217–237, ???? 1999. CODEN PAAPEC.

ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a778232861.

Pan:1996:EBT

[PN96] Zhengjun Pan and Lishan Nie. Evolving both the topology and weights of neural networks. Parallel Algorithms and Applications, 9(3–4):299–307, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a778707986.

Papadopoulou:1995:MFS

[PS95] Elena Papadopoulou and Yiannis Saridakis. Modular fixed-size VLSI architectures for general multisplitting iteration. Parallel Algorithms and Applications, 7(3–4):177–191, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772402405.

Pini:1999:PFE

[PS99] Giorgio Pini and Flavio Sartoretto. A parallel finite element code for nonlinear leaky aquifer systems. Parallel Algorithms and Applications, 13(4):345–361, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776633664.

Purcz:2003:NPA

[Pur03] Pavol Purcz. New parallel algorithm and its speed-up for the explicit difference method. Parallel Algorithms and Applications, 18(1–2):3–11, March/ June 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a716776277.

Algorithms and applications of parallel numerics (Bratislava, 2000).

Pinar:1993:CSP

[PZ93] Mustafa C. Pinar and Stavros A.

Zenios. A comparative study of parallel decompositions for multicommodity flow problems. Parallel Algorithms and Applications, 1(4):255–271, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397697.

Piccolomini:2003:EPS

[PZ03] E. Loli Piccolomini and F. Zama.

An experiment of parallel SPECT data reconstruction. Parallel Algorithms and Applications, 18(3):107–119, September 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713812562.

Qiu:1994:LBS

[QA94] K. Qiu and S. G. Akl. Load balancing, selection and sorting on the star and pancake interconnection networks. Parallel Algorithms and Applications, 2(1–2):27–42, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745440.

Qiming:1994:DDM

[QL94] He Qiming and Kang Lishan. The domain decomposition method for multidimensional and large scale nonlinear systems—case of subdomains overlapping. Parallel Algorithms and Applications, 4(1–2):77–90, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745101.

Qingping:2003:NCP

[QPPJ03] Guo Qingping, Yakup Paker, Dennis Parkinson, and Xiao JinSheng. Network computing performance evaluation in message passing environment. Parallel Algorithms and Applications, 18(4):217–224, December 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713646635.

Quintana:1995:EPS

[Qsn95a] Enrique S. Quintana, Mercedes Marqu s, and Vicente Hern ndez. An efficient parallel Sylvester equation solver based on the Hessenberg–Schur method. Parallel Algorithms and Applications, 7(1–2):133–141, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743297.

Quintana:1995:PTS

[Qsn95b] Enrique S. Quintana, Mercedes Marqu s, and Vicente Hern ndez. A parallel triangular Sylvester equation solver based on the Hessenberg–Schur method. Parallel Algorithms and Applications, 5(1–2):107–118, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743844.

Rantakokko:1998:FPS

[Ran98] Jarmo Rantakokko. A framework for partitioning structured grids with inhomogeneous workload. Parallel Algorithms and Applications, 13 (2):135–151, February 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776661630.

Rao:1994:PSL

[Rao94] S. Chandra Sekhara Rao. Parallel solution of the linear systems by an alternate quadrant interlocking factorization method. Parallel Algorithms and Applications, 4(1–2):1–20, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745096.

Rao:1996:EDU

[Rao96] Josyula R. Rao. Eventual determinism: Using probabilistic means to achieve deterministic ends. Parallel Algorithms and Applications, 8(1):3–19, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400357.

Rao:2000:PAB

[RDK00] S. Chandra Sekhara Rao, Pravir K. Dutt, and Mohan K. Kadalbajoo. A parallel algorithm for banded linear system. Parallel Algorithms and Applications, 14(3):235–252, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a778235469.

Redon:2000:DSP

[RF00] Xavier Redon and Paul Feautrier. Detection of scans in the polytope model. Parallel Algorithms and Applications, 15 (3–4):229–263, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776716069.

Renaut:1995:PMO

[RM95] R. A. Renaut and H. D. Mittelmann. Parallel multisplittings for optimization. Parallel Algorithms and Applications, 7(1–2):17–27, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743288.

Rajasekaran:1996:SSD

[RS96] S. Rajasekaran and S. Sahni. Sorting and selection on distributed memory bus computers. Parallel Algorithms and Applications, 8(3–4):179–193, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776632200.

Rajasethupathy:1994:TRS

[RTDL94] K. S. Rajasethupathy, Gwan-Ming Thio, S. K. Dhall, and S. Lakshmivarahan. Tridiagonalizing a real symmetric matrix: a parallel direct approach using Givens’ transformation. Parallel Algorithms and Applications, 2(4):305–313, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397563.

Sarbazi-Azad:2001:EAC

[SAOKM01] H. Sarbazi-Azad, M. Ould-Khaoua, and L. M. Mackenzie. Employing k-ary n-cubes for parallel Lagrange interpolation. Parallel Algorithms and Applications, 16(4):283–299, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777327099.

Sosonkina:2002:SAP

[SAW02] Masha Sosonkina, Donald C. S.

Allison, and Layne T. Watson. Scalability analysis of parallel GMRES implementations. Parallel Algorithms and Applications, 17(4):263–284, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713941500.

Stewart:2000:AFD

[SCG00] A. Stewart, M. Clint, and J. Gabarró. Axiomatic frameworks for developing BSP-style programs. Parallel Algorithms and Applications, 14(4):271–292, April 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778714134.

Schwandt:1995:PST

[Sch95] Hartmut Schwandt. The parallel solution of tridiagonal and block tridiagonal systems of equations by a Schur complement method. Parallel Algorithms and Applications, 6(2–3):207–227, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398626.

Schaffer:2000:CBM

[SCM00] Rainer Schaffer, Francky Catthoor, and Renate Merker. Combining background memory management and regular array copartitioning, illustrated on a full motion estimation kernel. Parallel Algorithms and Applications, 15(3–4):201–228, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776716068.

Stratulat:1995:VSM

[SE95] S. Stratulat and D. J. Evans. Virtual shared memory machines— an application of pvm. Parallel Algorithms and Applications, 7(1–2):143–160, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772743298.

Shen:1996:NMG

[SE96] Hong Shen and David J. Evans.

A new method for generating integer compositions in parallel. Parallel Algorithms and Applications, 9(1–2):101–109, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772744470.

Shen:1997:FTM

[SEY97] Hong Shen, David J. Evans, and Jia You. Fault-tolerant multicast with traffic-balancing in hypercubes. Parallel Algorithms and Applications, 11(3–4):287–298, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772742295.

Shen:2000:OMH

[She00] Hong Shen. Optimal multiselection in hypercubes. Parallel Algorithms and Applications, 14(3):203–212, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778235467.

Simos:1994:EPM

[Sim94] T. E. Simos. Efficient parallel methods for special second order ordinary differential equations. Parallel Algorithms and Applications, 2(4):245–250, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397558.

Song:1994:EHT

[SJL^{+}94] Chang G. Song, Jung-Sing Jwo, S. Lakshmivarahan, S. K. Dhall, John M. Lewis, and Christopher S. Velden. An experiment in hurricane track prediction using parallel computing methods. Parallel Algorithms and Applications, 2(4):315–332, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397564.

Svolos:2004:EBG

[SKK04] Andreas I. Svolos, Charalampos G. Konstantopoulos, and Christos Kaklamanis. Efficient binary and grey level morphological operations on a massively parallel processor. Parallel Algorithms and Applications, 19(1):33–50, March 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713646639.

Sklyarov:2002:HSM

[Skl02] V. Sklyarov. Hardware/software modeling of FPGA-based systems. Parallel Algorithms and Applications, 17(1):19–39, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777803176.

Suri:2001:DPI

[SM01] Neeraj Suri and Avi Mendelson. Design of a parallel interconnect based on communication pattern considerations.

Parallel Algorithms and Applications, 16(4):243–271, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777327097.

Symons:1994:PAP

[SNS94] Anthony Symons, V. Lakshmi Narasimhan, and Kurt Sterzl. Performance analysis of a parallel FFT algorithm on a transputer network. Parallel Algorithms and Applications, 4(1–2):111–124, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745103.

Sogno:1997:JTH

[Sog97] Jean-Claude Sogno. The Janus test: A hierarchical algorithm for computing direction and distance vectors. (extended version). Parallel Algorithms and Applications, 12(1–3):57–82, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a775978204.

Sahai:1995:PEB

[SP95] Akhil Sahai and L. M. Patnaik.

Performance evaluation of backpropagation neural network implementation on single program multiple data (SPMD) architecture. Parallel Algorithms and Applications, 7(3–4):267–281, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772402451.

Spencer:1994:PAM

[Spe94] Thomas H. Spencer. Parallel approximate matching. Parallel Algorithms and Applications, 2(1–2):115–121, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745445.

Sheu:1994:MAH

[SS94a] Jang-Ping Sheu and Ming-Yang Su. A multicast algorithm for hypercube multiprocessors. Parallel Algorithms and Applications, 2(4):277–290, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772397561.

Sinharoy:1994:FOW

[SS94b] Balaram Sinharoy and Boleslaw K. Szymanski. Finding optimum wavefront of parallel computation. Parallel Algorithms and Applications, 2(1–2):5–26, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772745439.

Sinharoy:1995:ASI

[SS95] Balaram Sinharoy and Boleslaw K. Szymanski. Announcement of a Special Issue of the Journal of Parallel Algorithms and Applications on Optimising Compilers for Parallel Languages. Parallel Algorithms and Applications, 7(3–4):313, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772402454.

Sinharoy:1997:ISI

[SS97a] Balaram Sinharoy and Boleslaw K.

Szymanski. Introduction: Special issue on optimising compilers for parallel languages. Parallel Algorithms and Applications, 12(1–3):1–4, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a775978201.

Sinharoy:1997:PCS

[SS97b] Balaram Sinharoy and Boleslaw K.

Szymanski. Parallelising compilers and systems. Parallel Algorithms and Applications, 12(1–3):5–20, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a775978202.

Shenoy:1999:ADP

[SSBK99] U. Nagaraj Shenoy, Y. N.

Srikant, V. P. Bhatkar, and Sandeep Kohli. Automatic data partitioning by hierarchical genetic search. Parallel Algorithms and Applications, 14(2):119–147, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777957923.

Schimmler:2004:BSF

[SSL04] Manfred Schimmler, Bertil Schmidt, and Hans-Werner Lang. A bit-serial floatingpoint unit for a massively parallel system on a chip. Parallel Algorithms and Applications, 19(2–3):79–95, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a714592649.

Stojmenovic:1993:RBI

[Sto93]  Ivan Stojmenović.  Routing

and broadcasting on incomplete and Gray code incomplete hypercubes. Parallel Algorithms and Applications, 1(3):167–177, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398090.

Stojmenovi:1994:ISI

[Sto94]  Ivan Stojmenovi.  Introduc-

tion to the Special Issue of Parallel Algorithms and Applications Devoted to the Parallel Algorithms Minitrack of HICSS-26. Parallel Algorithms and Applications, 2(1–2):1–4, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772745438.

Stoica:1996:TOA

[Sto96] Ion Stoica. Time-optimal algorithms for generalized dominance computation and related problems on mesh connected computers and meshes with multiple broadcasting. Parallel Algorithms and Applications, 8(3–4):271–283, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776632236.

Stpiczynski:2003:FPA

[Stp03] Przemysław Stpiczyński. Fast parallel algorithm for polynomial evaluation. Parallel Algorithms and Applications, 18 (4):209–216, December 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713646634.

Schurer:2003:EAN

[SU03] Rudolf Schürer and Andreas Uhl. An evaluation of adaptive numerical integration algorithms on parallel systems. Parallel Algorithms and Applications, 18(1–2):27–47, March/ June 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713812544.

Algorithms and applications of parallel numerics (Bratislava, 2000).

Sukon:1996:TPA

[Suk96] K. S. Sukon. On two parameter alternating group explicit (TAGE) method for singular perturbation problems. Parallel Algorithms and Applications, 10(1–2):71–77, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776220897.

Szularz:1999:RTL

[SWC99] M. Szularz, J. Weston, and M. Clint. Restarting techniques for the Lanczos algorithm and their implementation in parallel computing environments: Architectural influences. Parallel Algorithms and Applications, 14(1):57–77, ???? 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777980765.

Szularz:1995:MCL

[SWMC95] M. Szularz, J. Weston, K. Murphy, and M. Clint. Monitoring the convergence of the Lanczos algorithm in parallel computing environments. Parallel Algorithms and Applications, 6(4):287–302, September 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772403915.

Shyu:1998:EEX

[SWT98] Daw-Jong Shyu, Biing-Feng Wang, and Chuan-Yi Tang. Efficient emulations for Xtrees and m-ary trees. Parallel Algorithms and Applications, 13(2):95–116, February 1998. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776661628.

Talia:2001:MTP

[Tal01] Domenico Talia. Models and trends in parallel programming. Parallel Algorithms and Applications, 16(2–3):145–180, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777319807.

Turner:1996:TDL

[Td96] Gavin Turner and Heiko Schr der. Token distribution and load balancing on reconfigurable d-dimensional meshes. Parallel Algorithms and Applications, 8(3–4):251–269, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a776632235.

Tout:1993:DSD

[TE93] K. R. Tout and D. J. Evans.

Dynamic scheduled data-driven models for parallel expert systems. Parallel Algorithms and Applications, 1(4):285–301, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772397700.

Teng:1996:APA

[Ten96] S. Teng. Adaptative parallel algorithms for integral knapsack problem. Parallel Algorithms and Applications, 8(??):400–406, ???? 1996. CODEN PAAPEC. ISSN 1063-7192.

Tabirca:2002:PIF

[TFT02] T. Tabirca, L. Freeman, and S. Tabirca. On O(log p) parallel implementation of feedback guided dynamic loop scheduling. Parallel Algorithms and Applications, 17(2):157–164, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777924086.

Thune:1993:PAC

[Thu93]  Michael Thuné.  A parti-

tioning algorithm for composite grids. Parallel Algorithms and Applications, 1(1):69–81, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772399781.

Thune:1997:PSC

[Thu97]  Michael Thuné.  Partition-

ing strategies for composite grids. Parallel Algorithms and Applications, 11(3–4):325–348, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772742297.

Trobec:2003:NND

[TJ03] Roman Trobec and Uro Jovanovi. Node-to-node distances in regular d-neighbours topologies. Parallel Algorithms and Applications, 18(1–2):91–98, ???? 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a716776279.

Tsay:1995:NAF

[TK95] Jong-Chuang Tsay and Jenn-Yang Ke. A new approach to finding optimal linear schedules for uniform dependence algorithms. Parallel Algorithms and Applications, 7(1–2):73–86, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743292.

Tollenaere:1993:DML

[TOR93] Tom Tollenaere, Guy A. Orban, and Dirk Roose. Decomposition and mapping of locally connected layered neural networks on message-passing multiprocessors. Parallel Algorithms and Applications, 1(1):43–56, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772399756.

Taniar:2002:PSH

[TR02] David Taniar and J. Wenny Rahayu. Parallel sort-hash objectoriented collection join algorithms for shared-memory machines. Parallel Algorithms and Applications, 17(2):85–126, ???? 2002. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777924083.

Tout:1995:CPP

[TRFM95] K. Tout, J. Ribeiro-Filho, and B. Mignot. A cross-platform parallel genetic algorithms programming environment. Parallel Algorithms and Applications, 7(3–4):255–266, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772402450.

Tsay:1993:MTS

[Tsa93] Jyh-Jong Tsay. Mapping treestructured computations onto mesh-connected arrays of processors. Parallel Algorithms and Applications, 1(3):211–220, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772398093.

Thiele:2000:RSM

[TTS00] Lothar Thiele, Jürgen Teich, and Karsten Strehl. Regular state machines. Parallel Algorithms and Applications, 15(3–4):265–300, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776716064.

Uhl:1997:FIC

[UH97] A. Uhl and J. Hämmerle. Fractal image compression on MIMD architectures. I. Basic algorithms. Parallel Algorithms and Applications, 11(3–4):187–204, ???? 1997. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772742290.

Vyas:2003:COG

[VMG03] A. Vyas, R. Maloo, and P. Gupta. Cost-optimal Gaussian elimination algorithm on RCC–FULL. Parallel Algorithms and Applications, 18 (4):171–180, December 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713646631.

Vajtersic:2003:FSP

[VZ03] Marian Vajteršic and Peter Zinterhof. Foreword: Selected papers from the International Workshop “Parallel Numerics” held in Bratislava, September 18–20, 2000. Parallel Algorithms and Applications, 18(1–2):1–2, March/ June 2003. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a713812541.

Wang:1994:CAPa

[WBE94a] Deren Wang, Zhongzhi Bai, and D. J. Evans. A class of asynchronous parallel matrix multisplitting relaxation methods. Parallel Algorithms and Applications, 2(3):173–192, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401563.

Wang:1994:CAPb

[WBE94b] Deren Wang, Zhongzhi Bai, and D. J. Evans. A class of asynchronous parallel nonlinear multisplitting relaxation methods. Parallel Algorithms and Applications, 2(3):209–228, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401609.

Wan:1993:SAC

[WE93] C. R. Wan and D. J. Evans. A systolic architecture for Capon’s directions-of-arrival (doa) estimation method. Parallel Algorithms and Applications, 1(2):99–114, ???? 1993. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772401500.

Wu:1996:POB

[WH96] Po-Ting Wu and Elias N. Houstis. Parallel optimization based electronic prototyping of physical parts. Parallel Algorithms and Applications, 9(3–4):237–263, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a778707982.

Wilde:2000:LDP

[Wil00] Doran K. Wilde. A library for doing polyhedral operations. Parallel Algorithms and Applications, 15(3–4):137–166, ???? 2000. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776716066.

Wedler:2001:PIG

[WL01] Christoph Wedler and Christian Lengauer. On the parallel implementation of a generalized broadcast. Parallel Algorithms and Applications, 16(1):65–82, ???? 2001. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a777308747.

Williams:1994:CCI

[WMF94] Shirley A. Williams, Philip C. H. Mitchell, and Graham E. Fagg. A cluster computing implementation of a Monte Carlo simulation of a particle growth mechanism. Parallel Algorithms and Applications, 4(3–4):275–280, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a777314737.

Xing-Ping:1994:BVP

[XPJG94] Liu Xing-Ping and Hu Jia-Gan. Block vectorizable preconditioned iterative methods. Parallel Algorithms and Applications, 3(3–4):239–247, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779167059.

Xing-Ping:1996:MPI

[XPJG96] Liu Xing-Ping and Hu Jia-Gan. Multisplitting preconditioned iterative methods. Parallel Algorithms and Applications, 9(1–2):131–144, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772744473.

Yaakub:1995:RKC

[YE95a] A. R. Yaakub and D. J. Evans.

The Runge–Kutta contraharmonic mean method with extrapolation for the parallel solution of O.D.E.’s. Parallel Algorithms and Applications, 6(2–3):111–128, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772398621.

Yousif:1995:ECG

[YE95b] W. S. Yousif and D. J. Evans.

Explicit de-coupled group iterative methods and their parallel implementations. Parallel Algorithms and Applications, 7(1–2):53–71, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a772743291.

Yang:1999:SSL

[YL99] Tianruo Yang and Hai Xiang Lin. Solving sparse least squares problems with preconditioned CGLS method on parallel distributed memory computers. Parallel Algorithms and Applications, 13(4):289–305, April 1999. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a776633631.

Yong:1994:AEA

[YLE94] Liu Yong, Kang Lishan, and D. J. Evans. The annealing evolution algorithm. Parallel Algorithms and Applications, 3(1–2):145–150, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a779202348.

Yang:2004:DVC

[YM04] Xiaofan Yang and Graham M. Megson. On the double-vertex-cycle-connectivity of crossed cubes. Parallel Algorithms and Applications, 19(1):11–17, March 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a713646638.

Yuanxiang:1996:UMS

[YXZ96] Li Yuanxiang, Zou Xiufen, and Huang Zhangcan. Use of multi-scalar analysis for building lattice models of PDE’s. Parallel Algorithms and Applications, 9(3–4):265–272, ???? 1996. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a778707983.

Zhou:1994:EAD

[Zho94] Honbo B. Zhou. An effective approach for distributed program allocation. Parallel Algorithms and Applications, 3(1–2):57–71, ???? 1994. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a779202302.

Zama:1995:EIR

[ZP95] F. Zama and E. Loli Piccolomini. An experiment in image restoration using transputer networks. Parallel Algorithms and Applications, 5(3–4):283–292, ???? 1995. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a772400097.

Zafar:2004:DFD

[ZPBD04] Bilal Zafar, Timothy M. Pinkston, Aurelio Bermúdez, and Jos Duato. Deadlock-free dynamic reconfiguration over InfiniBand^{T M} networks. Parallel Algorithms and Applications, 19 (2–3):127–143, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

www.informaworld.com/smpp/
content~content=a714592651.

Zeyao:2004:AMI

[ZZ04] Mo Zeyao and Huang Zhengfeng. Application of MPI-IO in parallel particle transport Monte– Carlo simulation. Parallel Algorithms and Applications, 19(4):227–236, ???? 2004. CODEN PAAPEC. ISSN 1063-7192. URL http://

                                        www.informaworld.com/smpp/
                                        content~content=a714592658.